### v1

# cuVSLAM: CUDA accelerated visual odometry and mapping

Abstract
Accurate and robust pose estimation is a key requirement for any autonomous robot. We present cuVSLAM, a state-of-the-art solution for visual simultaneous localization and mapping, which can operate with a variety of visual-inertial sensor suites, including multiple RGB and depth cameras, and inertial measurement units. cuVSLAM supports operation with as few as one RGB camera to as many as 32 cameras, in arbitrary geometric configurations, thus supporting a wide range of robotic setups. cuVSLAM is specifically optimized using CUDA to deploy in real-time applications with minimal computational overhead on edge-computing devices such as the NVIDIA Jetson. We present the design and implementation of cuVSLAM, example use cases, and empirical results on several state-of-the-art benchmarks demonstrating the best-in-class performance of cuVSLAM.
[…]
Acknowledgements
We would like to extend our heartfelt acknowledgment to Eugene Vendrovskiy, Stanislav Volodarskiy, and Dmitry Robustov for their invaluable contributions to this project. Their profound expertise, exceptional skills, and commitment to excellence have been pivotal in shaping the outcomes of this work.
1 Introduction
Visual Simultaneous Localization and Mapping (VSLAM) is a fundamental capability for autonomous robots, enabling them to build a metric map of their environment while simultaneously determining their location within it. The increasing deployment of autonomous robots across diverse applications, from warehouse automation to last-mile delivery, has created a pressing need for robust, real-time VSLAM solutions that can operate efficiently across diverse settings, while being able to run on energy-efficient edge computing devices.
Traditional VSLAM systems face several key challenges: they often struggle with real-time performance on resource-constrained platforms, have limited flexibility in sensor configurations, and may not fully utilize available hardware acceleration capabilities. Additionally, many existing solutions are optimized for specific sensor configurations, making them less adaptable to diverse robotic platforms with varying sensor requirements.
We present cuVSLAM, a CUDA-accelerated visual SLAM system designed to address these challenges. cuVSLAM offers several key innovations:
  • Flexible sensor suite support, accommodating configurations from a single RGB camera to complex arrays of up to 32 cameras with arbitrary geometric arrangements
  • Support for optional IMU and depth sensors
  • Efficient CUDA implementation enabling real-time performance on edge devices like NVIDIA Jetson
  • A modular architecture separating frontend pose estimation from backend map refinement
  • Robust feature tracking and mapping capabilities that maintain accuracy across diverse environments
Our technical approach combines classical SLAM techniques with modern GPU acceleration. The frontend module performs real-time pose estimation using feature detection, tracking, and local mapping, while the backend handles global map consistency through pose graph optimization and loop closure. By leveraging CUDA acceleration throughout the pipeline, from feature detection to bundle adjustment, cuVSLAM achieves superior performance while maintaining high accuracy.
We evaluate cuVSLAM on standard benchmarks including KITTI, EuRoC, and TUM RGB-D, demonstrating state-of-the-art accuracy while maintaining real-time performance on edge computing platforms. Our results show that cuVSLAM achieves average trajectory errors below 1% on the KITTI odometry benchmark and mean position errors under 5cm on the EuRoC dataset, while processing frames in real-time on NVIDIA Jetson platforms.
The rest of this technical report is organized as follows: Section 2 describes the system architecture and design, Section 3 presents experimental results and benchmarks, and Section 4 concludes with discussion and future work.
2 System Architecture and Design
Figure 1: cuVSLAM architecture
The cuVSLAM library implements a general approach to camera pose estimation, where the architecture consists of two major blocks: the frontend, responsible for online pose estimation with minimal latency and maximum throughput, and the backend, focusing on asynchronous and potentially resource-demanding map refinement, loop closing, and pose graph optimization.
The design of the frontend pipeline prioritizes local pose estimation accuracy over global consistency, focusing on delivering smooth trajectories without the bumps introduced by loop closure mechanisms or PGO (Pose Graph Optimization). To achieve this, it maintains a local odometry map consisting of the last $N$ 6DOF camera poses (at keyframes) along with visible 3D landmarks and their observations. Besides the map, the frontend consists of the 2D module, responsible for feature selection, 2D feature tracking, and keyframe selection, and the 3D module, which performs pose estimation based on the local map, 2D information, and other data sources (e.g., IMU or depth measurements).
In contrast to the frontend, cuVSLAM’s backend focuses on global map and pose consistency. It processes frontend outputs and builds a globally consistent map that integrates camera poses, 2D observations, 3D landmarks, pose deltas, and visual features. The backend maintains global consistency in the map through asynchronous refinement using pose graph optimization and loop closing.
2.1 2D module
Figure 2: 2D module implements keyframe selection, feature selection and 2D tracking.
cuVSLAM implements a set of algorithms operating on the 2D level, which include keypoint selection and feature extraction, keypoint tracking, and keyframe selection.
The pipeline begins with keypoint selection, which is designed to identify high-contrast features and ensure an approximately uniform distribution of keypoints across the image. The image is first divided into non-overlapping patches, forming an $N\times M$ grid. In each patch, the algorithm selects the top $k$ keypoints based on the "Good Features to Track" measure (Shi and Tomasi (1993)), where $k$ is calculated such that the total number of keypoints across the entire image exceeds a predefined threshold, $K_{I}$.
$k>\left\lfloor\frac{K_{I}}{N*M}\right\rfloor$ (1)
For 2D tracking, cuVSLAM implements a modified version of the Lucas-Kanade algorithm (Lucas and Kanade (1981), Bouguet (2000)) for feature tracking that: 1) performs tracking in a coarse-to-fine manner, continuously refining track positions at each image pyramid level, and 2) performs a normalized cross-correlation (NCC) check at each optimization step for each keypoint to filter out unreliable tracks.
If the number of successfully tracked keypoints falls below a threshold, the 2D module creates a new keyframe, which triggers local and global map updates and asynchronous refinement. In stereo and multicamera modes, cuVSLAM launches cross-camera feature tracking to obtain multi-view observations for subsequent landmark triangulation.
2.2 3D module
When a keyframe is created, the 3D module collects multi-view observations from the 2D module and performs landmark triangulation. The created landmarks, along with their observations, are saved in the local map. After pose estimation, an asynchronous local sparse bundle adjustment (SBA) is initiated for refinement. Given the poses $T_{i}$, landmarks $p_{j}$, and 2D observations $o_{ij}$, SBA solves the following optimization task:
$r^{repr}_{ijk}=\pi(T^{cb}_{k}T^{bw}_{i}p^{w}_{j})-o_{j,k}$ (2)
$\hat{T}^{bw}_{1:N},\hat{p}^{w}_{1:M}=\arg_{T^{bw}_{1:N},p^{w}_{1:M}}\min\sum_{i\in[1,N]}\sum_{j\in[1,M]}\sum_{k\in[1,C]}||\pi(T^{cb}_{k}T^{bw}_{i}p^{w}_{j})-o_{j,k}||^{2}_{\Sigma}$ (3)
where $r^{repr}$ represents reprojection error, $T^{cb}_{k}$ refers to the k-th camera-from-base transformation, $T^{bw}_{i}$ denotes the base-from-world transformation, $p^{w}$ is the 3D landmark in the world frame, and $o_{j,k}$ represents the observation of landmark $j$ by the k-th camera.
Following the common approach, we use the Schur complement to solve for poses $T_{i}$ first and solve for points $p_{j}$ afterward. To enhance the efficiency of SBA further, we implement it in CUDA, which is particularly useful for edge devices and/or in multicamera setups. cuVSLAM supports mono, stereo, multicamera, and visual-inertial modes, each of which corresponds to a dedicated 3D solver.
2.2.1 Stereo
Stereo mode is the default mode in cuVSLAM. It utilizes PnP algorithm for pose estimation, that relies on $o_{j}$, 2D observation, and $p_{i}$, 3D landmarks, acquired from the local map.
$\hat{T}^{bw}=\arg_{T^{bw}}\min\sum_{j\in[1,M]}\sum_{k\in[1,2]}||r^{repr}_{jk}||^{2}_{\Sigma}$ (4)
On regular frames, we perform tracking only for the left camera for the tracks that are still active. We query the local map for the corresponding 3D landmarks and use them for pose estimation. When the number of successfully tracked keypoints falls below a threshold, a keyframe is created. After feature selection, we perform cross-camera (left-to-right) tracking to obtain landmark stereo observations and conduct point triangulation. Finally, we add the current camera pose, along with the created landmarks and their observations, into the local map and launch asynchronous SBA for refinement.
2.2.2 Multicamera
When a robot operates in environments where it is likely to face a featureless surface from one direction, e.g., narrow corridors or elevators, it becomes vital to use multiple cameras for pose estimation.
cuVSLAM extends the approach introduced in 2.2.1 to a multicamera setup by treating an arbitrary multiple-camera configuration as a number of stereo pairs.
At startup, we build a Frustum Intersection Graph (FIG)—a directed graph $(V,E)$, where vertices $V$ represent a set of optical sensors, and edges $E$ connect cameras that share a common field of view. Edge direction describes the cross-camera tracking between adjacent cameras, i.e., if $e_{ij}$ exists, cuVSLAM performs 2D tracking from camera $c_{i}$ to $c_{j}$.
The FIG is built automatically at cuVSLAM’s startup. Provided with a list of camera extrinsics/intrinsics, cuVSLAM iterates over each pair of cameras and creates a new edge in the graph when cameras have overlapping fields of view (FoV). To detect this, cuVSLAM places a virtual plane in front of the first camera. It selects uniformly distributed pixels in the image $I_{1}$, lifts them onto the plane, and backprojects them to the image $I_{2}$. When the ratio of successfully projected points to the total number of sampled points exceeds a threshold, cuVSLAM detects the shared FoV and creates an edge.
Figure 3: Randomly selected pixels on left picture are lifted on a virtual plane in front of the camera. Pixels that can be successfully projected onto second images are shown in green.
Cameras that correspond to the source of some edge are used for feature tracking/selection over time on each incoming frame. A keyframe is still treated as a global event, while feature selection happens independently for each camera.
Let $S_{curr}$ be a set of all currently observed points from all the cameras and $S_{kf}$ is the set of points observed in the previous keyframe. Then if:
$\frac{\left|S_{curr}\cap S_{kf}\right|}{\left|S_{kf}\right|}<T;$ (5)
a keyframe is created and cross-camera tracking is done. Similar to stereo mode we then perform landmark triangulation and refine the map with SBA.
After 2D tracking both on regular frames and keyframes, we collect all the available observations along with the corresponding 3D landmarks for pose estimation.
$\hat{T}^{bw}=\arg_{T^{bw}}\min\sum_{j\in[1,M]}\sum_{k\in[1,C]}||r^{repr}_{jk}||^{2}_{\Sigma}$ (6)
where $r^{repr}$ denotes a reprojection error for j-th landmark observed from k-th camera in the rig Barfoot (2017).
Since observations from multiple cameras contribute to pose estimation, cuVSLAM requires images to be synchronized. This is a minor restriction, as many modern cameras provide hardware synchronization capabilities, such as Intel RealSense RealSense_2018.
2.2.3 Visual-Inertial
Figure 4: Factor graphs for VIO pipeline.
Visual-Inertial (VI) mode is inspired by Forster et al. (2017) and Campos et al. (2021). Following their approach cuVSLAM defines robot‘s state as a combination of 6 DoF pose, linear velocity, accelerometer and gyroscope biases forming a 15 DoF state vector $S=[T\in SE(3),v\in\mathbb{R}^{3},b^{a}\in\mathbb{R}^{3},b^{w}\in\mathbb{R}^{3}]$. Robot states along with IMU ($r^{imu}$) and visual ($r^{repr}$) factors are used to build a factor-graph for a list of crucial subproblems, which are:
  • Purely visual pose estimation is used for the initialization of VI pipeline. It consists of the PnP and the asynchronous SBA as it was described in section 2.2.1 and needed to build a pose history for the subsequent gravity estimation.
  • Gravity estimation. Gravity is a key parameter needed for the IMU preintegration and IMU-aided pose estimation. Assuming the poses predicted by the visual-only part are fixed, the algorithm estimates velocities $v_{i}$, accelerometer and gyroscope biases $b^{a}$, $b^{w}$ and a rotation matrix $R^{g}$, s.t. $g=R^{g}[0,0,9.81]^{T}$.
    $R^{g},\textbf{b},v_{0:k}=argmin\left[\left\|b\right\|^{2}_{\Sigma_{prior}}+\sum_{i=0}^{k-1}\left\|r^{imu}(S_{i},S_{i+1})\right\|^{2}_{\Sigma_{IMU}}\right]$ (7)
  • Visual-inertial pose estimation is used every frame. It constrains two consecutive robot states $S_{i-1}$ and $S_{i}$ with IMU and visual factors. In addition $S_{i-1}$ is also constrained by the prior factor.
    $S_{i-1},S_{i}=argmin\left[\left\|r^{imu}(S_{i-1},S_{i})\right\|^{2}_{\Sigma_{IMU}}+\sum_{j=0}^{1}\left\|r^{repr}(S_{i-j})\right\|^{2}_{\Sigma_{vis}}+\left\|r^{prior}(S_{i-1})\right\|^{2}_{\Sigma_{p}}\right]$ (8)
  • Visual-inertial sparse bundle adjustment is used for local map refinement when IMU data along with gravity and bias estimations are available. Following prior works, we perform relinearization of the IMU factors based on the variable update norm or once per number of iterations.
    $S_{0:k}=argmin\left[\sum_{i=1}^{k}\left\|r^{imu}(S_{i-1},S_{i})\right\|^{2}_{\Sigma_{IMU}}+\sum_{i=0}^{k}\left\|r^{repr}(S_{i})\right\|^{2}_{\Sigma_{vis}}\right]$ (9)
2.2.4 RGBD
Figure 5: RGBD factor graph.
cuVSLAM supports mono RGBD-based dense frame-to-frame pose estimation. Following Kerl et al. (2013) it considers two consecutive intensity images $I_{1,2}$ along with pixel-aligned depth images $Z_{1,2}$ and estimates a relative 6 DoF transformation between corresponding frames.
It defines a list of factors, that constrain frame positions w.r.t each other:
  • Visual factor used to constrain poses w.r.t the local map by means of reprojection error $r^{repr}$ described earlier.
  • Dense intensity and depth factors define constraints for each pixel location $x$, introducing the following error functions:
    $r^{I}=\begin{bmatrix}I_{2}(\tau(T_{21},x))-I_{1}\\ I_{1}(\tau(T^{-1}_{21},x))-I_{2}\end{bmatrix}$ (10)
    $r^{Z}=\begin{bmatrix}Z_{2}(\tau(T_{21},x))-[T_{21}\pi^{-1}(x,Z_{1}(x))]_{z}\\ Z_{1}(\tau(T^{-1}_{21},x))-[T^{-1}_{21}\pi^{-1}(x,Z_{2}(x))]_{z}\end{bmatrix}$ (11)
    where $T_{21}$ is the relative pose; $\tau(T,x)$ is a warping function, that lifts the pixel location $x$ to 3D, applies transformation $T$ and projects the point back to image plane; [.$[.]_{z}$ returns Z coordinate of a point.
  • Point-to-point factor. Provided with the 2D-2D matches $(x,y)_{1:N}$ obtained in 2D module, we define the point-to-point error:
    $r^{p}=\pi^{-1}(y,Z_{2}(y))-T_{21}\pi^{-1}(x,Z_{1}(x))$ (12)
All together these factors form the combined optimization problem, which we solve using a Levenberg-Marquardt solver implemented on GPU:
$\hat{T_{21}}=\arg_{T_{21}}\min(\sum_{p\in Points}||r^{repr}(p)||^{2}_{\Sigma}+\sum_{x\in Pixels}(||r^{I}(x)||^{2}+||r^{Z}(x)||_{\sigma_{z}}^{2})+\sum_{(x,y)\in Tracks}||r^{p}(x,y)||^{2})$ (13)
2.2.5 Mono
Mono camera mode shares the same common approach as 2D feature tracking, with subsequent point triangulation and pose estimation. A key difference is that triangulation is performed using observations obtained from different camera poses, which we estimate using the same PnP algorithm as described in 2.2.1, given the current observations of previously obtained landmarks.
For the first frame pair, when no 3D landmarks are available, cuVSLAM estimates the relative transformation through the fundamental matrix. Specifically, given a set of 2D tracks $(x,y)_{1:N}$, we estimate the fundamental matrix $F$ using RANSAC. With the calibration matrix $K$, we then retrieve the up-to-scale relative pose.
The relative pose is subsequently used to triangulate the first set of landmarks and update the local map, followed by asynchronous SBA refinement.
2.3 Loop closing and global map refinement
Figure 6: Factor graphs for VIO pipeline.
cuVSLAM’s loop closing and global map refinement module helps limit global drift by enforcing global consistency of the map. It consists of 6 DoF keyframe poses, landmarks and their observations, pose deltas, and sparse image features. Pose deltas are obtained through the integration of intra-keyframe frontend pose predictions. For image features, we use a list of $9\times 9$ image patches taken from each level of the image pyramid.
We perform the following steps to detect loop closures and refine the map to enforce global consistency:
  • Integrate latest data. When a new keyframe is created, cuVSLAM extracts visual features, integrates intra-keyframe pose predictions, collects currently visible landmarks along with their observations and updates the global map using an asynchronous data queue.
  • Query nearest neighbors. After the map update, cuVSLAM performs a pose search of robot poses within a sphere of fixed radius. We use a kd-tree for fast and efficient spatial search.
  • Landmark selection. After collecting all the landmarks related to the found poses, we: 1) filter them by geometric visibility, 2) track landmark visual features onto the currently observed image, and 3) filter out landmarks with unsuccessful tracking. Finally, when landmarks from multiple poses are observed, we split the landmarks into sets with respect to connected poses and pick the largest set.
  • Relative pose estimation. At this step, we estimate a pose delta $D$ between the pose from the map and the current robot’s position. Let $T^{mw}$ denote a pose in the map, connecting the world and map frames. Then, after shifting landmarks into the map frame, we solve for the relative pose estimate:
    $\hat{T}^{bm}=\arg_{T^{bm}}\min\sum_{j\in[1,M]}\sum_{k\in[1,C]}||\pi(T^{cb}_{k}T^{bm}T^{mw}p^{w}_{j})-o_{j,k}||^{2}_{\Sigma}$ (14)
  • Pose graph optimization. Provided with the pose deltas between poses either by the frontend or the loop closure mechanism, cuVSLAM refines the map by minimizing the following sum across pose graph edges $E$:
    $T_{1:N}=\arg_{T_{1:N}}\min\sum_{i,j\in E}||Log(D_{ij}^{-1}T_{i}^{-1}T_{j})||^{2}$ (15)
3 Experiments
This section demonstrates the capabilities of the cuVSLAM library in robotics applications across diverse real-world environments and camera configurations. We also compare our solution with established classical and deep learning libraries that provide camera pose estimation for long-distance trajectories.
3.1 System Performance and Hardware Utilization
We evaluated cuVSLAM’s performance in real-time live execution to measure embedded hardware utilization, as well as using recordings from a warehouse environment to measure visual tracking execution time on embedded and workstation hardware. During these experiments, cuVSLAM processed feature-rich images captured either from a wheeled robot or from handheld camera movements, representing typical usage scenarios.
Table 1: cuVSLAM Performance Evaluation. The left side shows per-call processing times measured on RTX4090 (Desktop) and Orin AGX 64gb (Jetson), all images in datasets have 768×480 resolution (except mono-depth mode with resolution 640×480). The right side summarizes hardware utilization (CPU and GPU) measured on the Jetson platform running cuVSLAM via the Python API with live RealSense cameras at 640×480 resolution and 30 FPS
. Configuration Track call time, ms Jetson HW utilization Desktop Jetson CPU % GPU % Mono 0.9 2.7 NA NA Mono-Depth 5.9 13.9 2.4 45.0 Stereo 0.4 2.7 0.4 1.7 Stereo-Inertial 1.3 3.8 Multicamera (2-stereo) 0.8 2.7 1.1 5.9 Multicamera (3-stereo) 1.4 2.8 1.5 7.1 Multicamera (4-stereo) 2.1 2.8 NA NA
The left side of Tab. 1 summarizes execution times for the cuVSLAM visual tracking function, excluding image acquisition and result post-processing overhead. Performance was measured on two hardware configurations: a desktop workstation equipped with an Intel i7-14700 CPU and NVIDIA RTX 4090 GPU, and an embedded single-board Jetson Orin AGX 64GB platform running in MAXN power mode.
The right side of Tab. 1 presents CPU and GPU utilization percentages for cuVSLAM visual odometry executed through the Python API with RealSense D435/D455 cameras running at 30 FPS. These measurements were conducted live, in real-time. To isolate cuVSLAM’s hardware resource requirements, we performed two controlled experiments: first, we measured baseline system utilization with the RealSense camera in motion but without cuVSLAM processing input data; then, we repeated the identical camera movement pattern with cuVSLAM visual tracking enabled. Resource utilization metrics were sampled at 15 ms intervals for identical durations in both experiments. cuVSLAM’s net resource consumption was calculated by subtracting baseline measurements from total utilization and averaging the results. This approach accurately reflects the additional computational overhead introduced by the library. Note that these measurements were performed in a controlled environment, excluding other resource-intensive operations such as visualization, complex image processing, or data recording, which could otherwise skew results by overloading the system.
3.2 Benchmarking
Table 2: Evaluation results for cuVSLAM Visual Odometry and SLAM across various operating modes
Mode Dataset Method ATE, % ARE, deg APE (RMSE)
Mono-Depth ICL-Nuim ORBSLAM3 - - -
DPVO 10.28 0.77 0.61
cuVSLAM Odom 0.41 0.99 0.026
cuVSLAM SLAM 0.44 0.97 0.026
TUM RGB-D ORBSLAM3 0.50 2.98 0.067
DPVO 13.56 1.98 0.80
cuVSLAM Odom 1.35 5.52 0.11
cuVSLAM SLAM 0.99 4.13 0.065
Stereo EuroC* ORBSLAM3 0.21 1.41 0.068
DPVO 0.21 0.96 0.10
cuVSLAM Odom 0.29 1.96 0.13
cuVSLAM SLAM 0.17 1.12 0.054
Kitti ORBSLAM3 0.31 1.20 2.98
DPVO 21.69 1.14 195.05
cuVSLAM Odom 0.33 1.14 3.00
cuVSLAM SLAM 0.27 0.93 1.98
TartanAir ORBSLAM3 - - -
DPVO 12.73 0.79 12.84
cuVSLAM Odom 0.44 1.78 0.44
cuVSLAM SLAM 0.49 1.85 0.50
Stereo-Inertial EuroC ORBSLAM3 5.70 72.23 0.066
DPVO - - -
cuVSLAM Odom 0.39 2.69 0.19
cuVSLAM SLAM 0.29 2.27 0.13
Multi-Stereo R2B ORBSLAM3 - - -
DPVO 5.35 0.81 12.30
cuVSLAM Odom 0.18 1.15 0.28
cuVSLAM SLAM 0.11 0.70 0.18
The cuVSLAM operation modes were evaluated using several publicly available datasets: Mono-Depth mode was validated on the ICL-NUIM Handa_2014 and TUM RGB-D Sturm_2012 datasets; Stereo mode was tested on the EuroC Burri_2016, KITTI Geiger_2013, and TartanAir Wang_2020 datasets; and Stereo-Inertial mode was assessed using the EuroC dataset. Standard metrics were employed to quantify the accuracy of the estimated trajectories: Average Relative Translation and Rotation Errors proposed in Geiger_2013, as well as Absolute Pose Error with trajectory alignment and scale correction Sturm_2012 using the EVO library Grupp_2017.
To ensure accurate and representative comparisons, clearly identified outlier sequences were excluded. For instance, the EuroC V203 sequence was included only in Stereo-Inertial evaluations, and certain sequences from the TartanAir and TUM RGB-D datasets were omitted. A complete list of the sequences used from each dataset is provided in Appendix A.
To rigorously validate cuVSLAM’s multi-camera capabilities, we collected a proprietary Multi-Stereo R2B dataset from office and warehouse environments using the Nova Carter wheeled robot equipped with four stereo cameras. Example scenes are shown in Fig. 7, and trajectories are illustrated in Fig. 10. A detailed description of this dataset is presented in Appendix B. Ground truth trajectories for these evaluations were obtained using high-precision lidar localization.
The use of a proprietary dataset was necessary because cuVSLAM’s multi-camera mode requires images captured by hardware-synchronized cameras, which could be combined in multiple stereo pairs. Currently, publicly available multi-camera datasets generally lack either hardware-synchronized cameras suitable for stereo pair arrangements NCTL_2015, Hilti_2024 or employ software synchronization that does not ensure simultaneous image capture M2dgr_2021, CODa_2023.
We anticipate that recent growth in availability and adoption of synchronized multi-stereo camera systems from various vendors will encourage creation of new public multi-camera datasets, facilitating further development, testing, and implementation of multi-camera SLAM systems in robotic applications.
Figure 7: Sample camera views from the R2B dataset captured across diverse environments including large office spaces, industrial warehouses, and narrow corridors in both office and storage areas. The images show simultaneous captures from 4 stereo cameras (displaying only the left image from each stereo pair) taken by the Carter Robot
Tab. 2 presents comprehensive benchmarking results across popular datasets from various domains. For comparative analysis, we include performance metrics for the classical computer vision-based ORB-SLAM3 and the deep learning-based DPVO alongside cuVSLAM. Each library was configured appropriately for the dataset domain: ORB-SLAM3 and cuVSLAM operated in their corresponding specialized modes (monocular-depth, stereo, or inertial), while DPVO was consistently used in monocular mode across all evaluations (utilizing the front left camera for stereo and multi-camera datasets).
4 Discussion
cuVSLAM demonstrates robust performance across diverse robotic platforms, including automotive vehicles, aerial drones, wheeled robots, and handheld camera setups. Our evaluation covers both real-world and synthetic datasets, providing a comprehensive assessment of the system’s capabilities under varied operational conditions.
The presented results reflect cuVSLAM’s out-of-the-box performance without parameter fine-tuning, relying solely on standard sensor calibration parameters and stable image acquisition. This approach enables successful operation with various camera types and leaves room for potential improvements by end users. For example, in our TUM-RGBD dataset evaluation, raw depth data were used without any preprocessing, despite known limitations in depth quality for this dataset.
4.1 Hardware Scalability and Implementation
cuVSLAM exhibits excellent scalability across hardware platforms and provides effective integration through user-friendly ROS2111https://github.com/NVIDIA-ISAAC-ROS/isaac_ros_visual_slam and Python [^2] https://github.com/NVlabs/PyCuVSLAM APIs. Although this paper primarily considers scenarios involving multiple USB-connected RealSense cameras operating at 30 FPS, the measured visual tracking execution times and hardware utilization indicate significant potential for more demanding configurations. This is confirmed by practical experience, as cuVSLAM has been successfully deployed to process 8 Full HD distorted RGB images at 30 FPS from 4 stereo cameras in real-time on a Jetson Orin AGX as part of the ISAAC Perceptor framework [^3] https://developer.nvidia.com/isaac/perceptor, which simultaneously runs other resource-intensive components, such as NVBLOX Millane_2023 and Nav2 Macenski_2020.
4.2 Advantages and Research Directions
We advocate for wider adoption of multi-camera visual tracking in robotics based on two principal observed advantages: enhanced trajectory reliability in feature-poor environments in odometry mode, and increased loop closure detection rates in SLAM mode. Fig. 8 demonstrates trajectory reliability through a practical test conducted using a Carter robot equipped with four stereo cameras. In this experiment, the robot moved continuously while its cameras were randomly covered with non-transparent film for intervals of 20–60 seconds, ensuring at least one stereo pair remained uncovered at given time. Despite these challenging conditions, the odometry trajectory remained stable without jumps or discontinuities, validating the system’s robustness to partial visual obstruction.
Figure 8: Occlusion resilience test results using the Carter robot. The figure shows the continuous trajectory successfully maintained by cuVSLAM multicamera visual odometry despite random camera occlusions using non-transparent film. Numbered image rows correspond to instantaneous views from the robot’s four stereo cameras (showing only the left image from each stereo pair) at the respective numbered locations along the trajectory.
Additionally, the increased loop closure detection rate resulting from an expanded collective field of view is demonstrated in Fig. 9, comparing the number of loop closure events between single-stereo and four-stereo camera setups for the same trajectory. This higher frequency of camera pose adjustments yields a quantitative improvement of approximately 40% in SLAM accuracy metrics compared to odometry mode, as demonstrated in Tab. 2 using the Multi-Stereo R2B dataset.
Figure 9: Comparison of loop closure detection events between single-camera and multi-camera (four stereo pairs) SLAM configurations on identical trajectories. The multi-camera setup demonstrates improved mapping consistency due to its wider cumulative field-of-view.
Appendix A Appendix
A.1 Appendix A
The following publicly available datasets were used for validating the libraries:
The TartanAir dataset includes 120 Hard sequences, the complete list is provided in Tab. 3 . From the TUM RGB-D SLAM dataset, only the sequences from the freiburg3 subset were utilized, as these sequences specifically provide rectified images required by the cuVSLAM Mono-Depth mode. The 10 sequences employed from this subset are listed in Tab. 4.
Table 3: TartanAir dataset sequences used for evaluation
Environment Sequences
abandonedfactory P000-P011
amusement P000-P007
carwelding P000-P003
endofworld P000-P002, P005-P006
hospital P037-P039, P042, P044-P049
japanesealley P000-P005
neighborhood P000-P002, P004,
P006-P010, P012-P017
ocean P000-P008
office2 P000-P002, P004-P008, P010
office P000, P001, P004, P006
oldtown P000-P008
seasidetown P000, P002, P004
seasonsforest P001-P002, P004-P006
seasonsforest_winter P010-P018
soulcity P000-P002, P004, P008, P009
westerndesert P000, P002, P003, P005-P007
Table 4: TUM RGB-D SLAM dataset sequences used for evaluation
Sequences
large cabinet validation
long office household
nostructure texture far
nostructure texture near withloop
sitting halfsphere
sitting xyz
sitting xyz validation
structure texture far
structure texture near
teddy
A.2 Appendix B
The Multi-Stereo R2B dataset was collected using a wheeled Carter robot equipped with four Full HD RGB stereo Hawk cameras operating at 30 FPS. Stereo cameras are hardware-synchronized, with a maximum synchronization time difference of approximately 0.1 ms within each 8-image batch. Each recording has been validated for completeness regarding lost image batches; the maximum loss rate of image batches is below 0.04%. Fig. 10 shows the trajectories of all sequences along with detailed statistics, including the trajectory distance and the number of synchronized image batches for each recording.
Figure 10: Trajectories of sequences from the Multi-Stereo R2B dataset, along with corresponding distances and the number of 8-image batches per trajectory

### v2

@@
 Additionally, many existing solutions are optimized for specific sensor configurations, making them less adaptable to diverse robotic platforms with varying sensor requirements.
+We evaluate cuVSLAM on standard benchmarks, demonstrating state-of-the-art accuracy while maintaining real-time performance on edge computing platforms.
+Our results show that cuVSLAM achieves average trajectory errors below 1% on the KITTI odometry benchmark and position errors under 5cm on the EuRoC dataset, while processing frames in real-time on NVIDIA Jetson platforms.
+Multi-Stereo mode significantly improves accuracy and robustness on challenging sequences compared to single stereo cameras, as validated through both simulated and real-world dataset evaluations.
 We present cuVSLAM, a CUDA-accelerated visual SLAM system designed to address these challenges. cuVSLAM offers several key innovations:
@@
 By leveraging CUDA acceleration throughout the pipeline, from feature detection to bundle adjustment, cuVSLAM achieves superior performance while maintaining high accuracy.
-We evaluate cuVSLAM on standard benchmarks including KITTI, EuRoC, and TUM RGB-D, demonstrating state-of-the-art accuracy while maintaining real-time performance on edge computing platforms.
-Our results show that cuVSLAM achieves average trajectory errors below 1% on the KITTI odometry benchmark and mean position errors under 5cm on the EuRoC dataset, while processing frames in real-time on NVIDIA Jetson platforms.
 The rest of this technical report is organized as follows: Section 2 describes the system architecture and design, Section 3 presents experimental results and benchmarks, and Section 4 concludes with discussion and future work.
@@
 Finally, we add the current camera pose, along with the created landmarks and their observations, into the local map and launch asynchronous SBA for refinement.
-2.2.2 Multicamera
+2.2.2 Multi-Stereo
 When a robot operates in environments where it is likely to face a featureless surface from one direction, e.g., narrow corridors or elevators, it becomes vital to use multiple cameras for pose estimation.
@@
 Since observations from multiple cameras contribute to pose estimation, cuVSLAM requires images to be synchronized.
-This is a minor restriction, as many modern cameras provide hardware synchronization capabilities, such as Intel RealSense RealSense_2018.
+This is a minor restriction, as many modern cameras provide hardware synchronization capabilities, such as Intel RealSense Grunnet-Jepsen et al.
+(2018).
 2.2.3 Visual-Inertial
@@
     $S_{0:k}=argmin\left[\sum_{i=1}^{k}\left\|r^{imu}(S_{i-1},S_{i})\right\|^{2}_{\Sigma_{IMU}}+\sum_{i=0}^{k}\left\|r^{repr}(S_{i})\right\|^{2}_{\Sigma_{vis}}\right]$ (9)
-2.2.4 RGBD
+2.2.4 Mono-Depth
 Figure 5: RGBD factor graph.
@@
 3.1 System Performance and Hardware Utilization
-We evaluated cuVSLAM’s performance in real-time live execution to measure embedded hardware utilization, as well as using recordings from a warehouse environment to measure visual tracking execution time on embedded and workstation hardware.
-During these experiments, cuVSLAM processed feature-rich images captured either from a wheeled robot or from handheld camera movements, representing typical usage scenarios.
+We evaluated cuVSLAM’s performance through two complementary approaches: real-time execution on embedded hardware to measure resource utilization and single-frame processing time, and offline processing of warehouse environment recordings on workstation hardware.
+The test scenarios involved processing feature-rich images captured from either a wheeled robot navigating a warehouse environment or handheld camera movements.
+To ensure accurate measurements, all experiments were conducted on systems dedicated solely to cuVSLAM execution, without concurrent resource-intensive operations such as visualization, complex image processing, or data recording that could potentially skew the results.
+Table 1 summarizes the key performance metrics, with comprehensive results provided in Appendix A.2.
+The left portion of Table 1 presents execution times for cuVSLAM’s visual tracking function, excluding image acquisition and post-processing overhead.
+These measurements were obtained using recorded color images at 768×480 resolution (with the exception of mono-depth mode at 640×480 resolution) on two distinct hardware configurations: a desktop workstation featuring an Intel i7-14700 CPU with NVIDIA RTX 4090 GPU, and an embedded Jetson Orin AGX 64GB single-board computer operating in MAXN power mode.
 Table 1: cuVSLAM Performance Evaluation.
-The left side shows per-call processing times measured on RTX4090 (Desktop) and Orin AGX 64gb (Jetson), all images in datasets have 768×480 resolution (except mono-depth mode with resolution 640×480).
-The right side summarizes hardware utilization (CPU and GPU) measured on the Jetson platform running cuVSLAM via the Python API with live RealSense cameras at 640×480 resolution and 30 FPS
+Left: Per-call processing times measured on RTX 4090 (Desktop) and Jetson AGX Orin (64 GB) using datasets with image resolution 768×480 (except mono-depth mode at 640×480).
+Right: CPU and GPU utilization on Jetson AGX Orin during live operation with RealSense cameras at 640×480 resolution and 60 FPS (except stereo-inertial mode at 30 FPS).
 .
-Configuration Track call time, ms Jetson HW utilization Desktop Jetson CPU % GPU % Mono 0.9 2.7 NA NA Mono-Depth 5.9 13.9 2.4 45.0 Stereo 0.4 2.7 0.4 1.7 Stereo-Inertial 1.3 3.8 Multicamera (2-stereo) 0.8 2.7 1.1 5.9 Multicamera (3-stereo) 1.4 2.8 1.5 7.1 Multicamera (4-stereo) 2.1 2.8 NA NA
-The left side of Tab.
-1 summarizes execution times for the cuVSLAM visual tracking function, excluding image acquisition and result post-processing overhead.
-Performance was measured on two hardware configurations: a desktop workstation equipped with an Intel i7-14700 CPU and NVIDIA RTX 4090 GPU, and an embedded single-board Jetson Orin AGX 64GB platform running in MAXN power mode.
-The right side of Tab.
-1 presents CPU and GPU utilization percentages for cuVSLAM visual odometry executed through the Python API with RealSense D435/D455 cameras running at 30 FPS.
-These measurements were conducted live, in real-time.
-To isolate cuVSLAM’s hardware resource requirements, we performed two controlled experiments: first, we measured baseline system utilization with the RealSense camera in motion but without cuVSLAM processing input data; then, we repeated the identical camera movement pattern with cuVSLAM visual tracking enabled.
-Resource utilization metrics were sampled at 15 ms intervals for identical durations in both experiments.
-cuVSLAM’s net resource consumption was calculated by subtracting baseline measurements from total utilization and averaging the results.
-This approach accurately reflects the additional computational overhead introduced by the library.
-Note that these measurements were performed in a controlled environment, excluding other resource-intensive operations such as visualization, complex image processing, or data recording, which could otherwise skew results by overloading the system.
+Configuration Track call time, ms Jetson HW utilization Desktop Jetson CPU % GPU % Mono 0.9 2.7 NA NA Mono-Depth 5.9 15.1 2.6 55.0 Stereo-Inertial 1.3 3.8 1.3 2.2 Stereo 0.4 1.8 5.5 1.7 Multicamera (2-stereo) 0.8 2.0 8.3 9.0 Multicamera (3-stereo) 1.4 2.3 8.3 11.0 Multicamera (4-stereo) 2.1 NA NA NA
+The right portion of Table 1 details CPU and GPU utilization percentages for cuVSLAM visual odometry executed through ISAAC ROS 3.2 with RealSense D435/D455 stereo cameras.
+All live tests utilized cameras configured at 640×480 resolution and 60 FPS, except for Stereo-Inertial mode, which ran at 30 FPS with IMU data at 200 Hz.
+The reduced framerate was necessary to ensure valid IMU-based integration, providing 6-7 IMU measurements between each stereo image pair.
+To accurately isolate cuVSLAM’s hardware requirements, we employed a controlled experimental methodology: we first established baseline system utilization with the RealSense camera in motion but without cuVSLAM processing, then repeated the similar camera movement pattern with cuVSLAM visual tracking active.
+Resource utilization metrics were sampled at 15 ms intervals for equal durations in both experiments.
+The net resource consumption attributable to cuVSLAM was calculated by subtracting baseline measurements from total utilization values and averaging the results.
+This methodology provides an accurate assessment of the computational overhead introduced by the library.
+Additional evaluation details, including performance benchmarks for cuVSLAM executed via Python API on entry-level Jetson Orin Nano hardware, are presented in Appendix A.2.
 3.2 Benchmarking
+We evaluated cuVSLAM’s different operation modes using several publicly available datasets.
+The Mono-Depth mode was validated on the ICL-NUIM Handa et al.
+(2014) and TUM RGB-D Sturm et al.
+(2012) datasets.
+The Stereo mode was tested on the EuRoC Burri et al.
+(2016) and KITTI Geiger et al.
+(2013) datasets.
+The Stereo-Inertial mode was assessed using the EuRoC dataset.
+The Multi-Stereo mode was evaluated on simulated TartanAir V2 Wang et al.
+(2023) and TartanGround Patel et al.
+(2025) datasets and our proprietary R2B real-world dataset.
+To quantify trajectory estimation accuracy, we employed standard metrics: Average Relative Translation and Rotation Errors Geiger et al.
+(2013), and Absolute Pose Error with trajectory alignment and scale correction Sturm et al.
+(2012) using the EVO library Grupp (2017).
 Table 2: Evaluation results for cuVSLAM Visual Odometry and SLAM across various operating modes
@@
 cuVSLAM SLAM 0.27 0.93 1.98
-TartanAir ORBSLAM3 - - -
-DPVO 12.73 0.79 12.84
-cuVSLAM Odom 0.44 1.78 0.44
-cuVSLAM SLAM 0.49 1.85 0.50
 Stereo-Inertial EuroC ORBSLAM3 5.70 72.23 0.066
@@
 cuVSLAM SLAM 0.29 2.27 0.13
-Multi-Stereo R2B ORBSLAM3 - - -
-DPVO 5.35 0.81 12.30
-cuVSLAM Odom 0.18 1.15 0.28
+Table 2 presents comprehensive benchmarking results across popular datasets from various domains.
+For comparative analysis, we include performance metrics for the classical computer vision-based ORB-SLAM3 and the deep learning-based DPVO alongside cuVSLAM.
+Each library was configured according to the dataset characteristics: ORB-SLAM3 and cuVSLAM operated in their corresponding specialized modes (monocular-depth, stereo, or stereo-inertial), while DPVO consistently used monocular mode across all evaluations, utilizing the front-left camera for stereo datasets.
+Since scale corrections were applied to final trajectories rather than successive pose estimations, translation results may not be meaningful for some datasets; therefore, relative rotation error should be considered the primary metric.
+To ensure accurate and representative comparisons, we excluded clearly identified outlier and corrupted sequences.
+For instance, the EuRoC V203 sequence was included only in Stereo-Inertial evaluations, and only a subset of TUM RGB-D sequences were used (see Table 4).
+For TartanGround datasets, not all sequences provide the full complement of cameras or are available for download.
+Additionally, many sequences lack overlapping trajectories, preventing loop closure detection—a critical component for evaluating SLAM performance versus odometry.
+Consequently, we validated SLAM mode on selected subsets of TartanGround (see Table 7) and TartanAir V2 (see Tables 9 and 9) datasets.
+Figure 7: Sample camera views from the R2B dataset captured across diverse environments including large office spaces, industrial warehouses, and narrow corridors in both office and storage areas.
+The images show simultaneous captures from 4 stereo cameras (displaying only the left image from each stereo pair).
+While the new Tartan datasets provide free camera and quadruped robot movements in diverse simulated environments, we needed to rigorously validate cuVSLAM’s multi-camera capabilities in real-world scenarios.
+We therefore collected a proprietary Multi-Stereo R2B dataset in office and warehouse environments using the Nova Carter wheeled robot equipped with four stereo cameras.
+The example views from the R2B dataset are shown in Fig.
+7.
+This was our only available option for validating this cuVSLAM mode on real data, as publicly available real-world multi-camera datasets either lack hardware-synchronized cameras suitable for stereo pair arrangements Carlevaris-Bianco et al.
+(2015); Nair et al.
+(2024) or employ software synchronization that cannot ensure simultaneous image capture Yin et al.
+(2021); Zhang et al.
+(2023).
+Table 3 presents cuVSLAM’s results in Multi-Stereo Camera Mode.
+Notably, results achieved on the real R2B dataset closely match those from similar domains in the TartanGround dataset, confirming the validity of both datasets and results.
+For TartanAir V2, we performed detailed validation only on the Hard dataset, which presents greater challenges due to fast six-degrees-of-freedom motion and higher translational and angular velocities.
+Results for the Easy mode are comparable to those for TartanGround, with a detailed comparison available in Table 6.
+While TartanAir V2 results indicate room for improvement, they also demonstrate the importance of multi-camera mode for robust tracking in most environments, with some showing 2–4× improvement in multi-camera SLAM mode.
+Table 3: Evaluation results for cuVSLAM Multi-Stereo Visual Odometry and SLAM
+Dataset Method ATE, % ARE, deg APE (RMSE)
+R2B cuVSLAM Odom 0.18 1.15 0.28
 cuVSLAM SLAM 0.11 0.70 0.18
-The cuVSLAM operation modes were evaluated using several publicly available datasets: Mono-Depth mode was validated on the ICL-NUIM Handa_2014 and TUM RGB-D Sturm_2012 datasets; Stereo mode was tested on the EuroC Burri_2016, KITTI Geiger_2013, and TartanAir Wang_2020 datasets; and Stereo-Inertial mode was assessed using the EuroC dataset.
-Standard metrics were employed to quantify the accuracy of the estimated trajectories: Average Relative Translation and Rotation Errors proposed in Geiger_2013, as well as Absolute Pose Error with trajectory alignment and scale correction Sturm_2012 using the EVO library Grupp_2017.
-To ensure accurate and representative comparisons, clearly identified outlier sequences were excluded.
-For instance, the EuroC V203 sequence was included only in Stereo-Inertial evaluations, and certain sequences from the TartanAir and TUM RGB-D datasets were omitted.
-A complete list of the sequences used from each dataset is provided in Appendix A.
-To rigorously validate cuVSLAM’s multi-camera capabilities, we collected a proprietary Multi-Stereo R2B dataset from office and warehouse environments using the Nova Carter wheeled robot equipped with four stereo cameras.
-Example scenes are shown in Fig.
-7, and trajectories are illustrated in Fig.
-10.
-A detailed description of this dataset is presented in Appendix B.
-Ground truth trajectories for these evaluations were obtained using high-precision lidar localization.
-The use of a proprietary dataset was necessary because cuVSLAM’s multi-camera mode requires images captured by hardware-synchronized cameras, which could be combined in multiple stereo pairs.
-Currently, publicly available multi-camera datasets generally lack either hardware-synchronized cameras suitable for stereo pair arrangements NCTL_2015, Hilti_2024 or employ software synchronization that does not ensure simultaneous image capture M2dgr_2021, CODa_2023.
-We anticipate that recent growth in availability and adoption of synchronized multi-stereo camera systems from various vendors will encourage creation of new public multi-camera datasets, facilitating further development, testing, and implementation of multi-camera SLAM systems in robotic applications.
-Figure 7: Sample camera views from the R2B dataset captured across diverse environments including large office spaces, industrial warehouses, and narrow corridors in both office and storage areas.
-The images show simultaneous captures from 4 stereo cameras (displaying only the left image from each stereo pair) taken by the Carter Robot
-Tab.
-2 presents comprehensive benchmarking results across popular datasets from various domains.
-For comparative analysis, we include performance metrics for the classical computer vision-based ORB-SLAM3 and the deep learning-based DPVO alongside cuVSLAM.
-Each library was configured appropriately for the dataset domain: ORB-SLAM3 and cuVSLAM operated in their corresponding specialized modes (monocular-depth, stereo, or inertial), while DPVO was consistently used in monocular mode across all evaluations (utilizing the front left camera for stereo and multi-camera datasets).
+TartanGround cuVSLAM Odom* 0.54 0.95 0.14
+cuVSLAM Odom 0.21 0.48 0.09
+cuVSLAM SLAM 0.17 0.37 0.07
+TartanAir V2 (Hard) cuVSLAM Odom* 2.56 13.97 5.20
+cuVSLAM Odom 2.44 13.98 5.24
+cuVSLAM SLAM 2.26 12.76 4.99
+We anticipate that the recent growth in availability and adoption of synchronized multi-stereo camera systems from various vendors will encourage the creation of new public multi-camera datasets, facilitating further development, testing, and implementation of multi-camera SLAM systems in robotic applications.
 4 Discussion
-cuVSLAM demonstrates robust performance across diverse robotic platforms, including automotive vehicles, aerial drones, wheeled robots, and handheld camera setups.
-Our evaluation covers both real-world and synthetic datasets, providing a comprehensive assessment of the system’s capabilities under varied operational conditions.
-The presented results reflect cuVSLAM’s out-of-the-box performance without parameter fine-tuning, relying solely on standard sensor calibration parameters and stable image acquisition.
-This approach enables successful operation with various camera types and leaves room for potential improvements by end users.
-For example, in our TUM-RGBD dataset evaluation, raw depth data were used without any preprocessing, despite known limitations in depth quality for this dataset.
-4.1 Hardware Scalability and Implementation
-cuVSLAM exhibits excellent scalability across hardware platforms and provides effective integration through user-friendly ROS2111https://github.com/NVIDIA-ISAAC-ROS/isaac_ros_visual_slam and Python [^2] https://github.com/NVlabs/PyCuVSLAM APIs.
-Although this paper primarily considers scenarios involving multiple USB-connected RealSense cameras operating at 30 FPS, the measured visual tracking execution times and hardware utilization indicate significant potential for more demanding configurations.
-This is confirmed by practical experience, as cuVSLAM has been successfully deployed to process 8 Full HD distorted RGB images at 30 FPS from 4 stereo cameras in real-time on a Jetson Orin AGX as part of the ISAAC Perceptor framework [^3] https://developer.nvidia.com/isaac/perceptor, which simultaneously runs other resource-intensive components, such as NVBLOX Millane_2023 and Nav2 Macenski_2020.
-4.2 Advantages and Research Directions
-We advocate for wider adoption of multi-camera visual tracking in robotics based on two principal observed advantages: enhanced trajectory reliability in feature-poor environments in odometry mode, and increased loop closure detection rates in SLAM mode.
-Fig.
-8 demonstrates trajectory reliability through a practical test conducted using a Carter robot equipped with four stereo cameras.
-In this experiment, the robot moved continuously while its cameras were randomly covered with non-transparent film for intervals of 20–60 seconds, ensuring at least one stereo pair remained uncovered at given time.
-Despite these challenging conditions, the odometry trajectory remained stable without jumps or discontinuities, validating the system’s robustness to partial visual obstruction.
+We advocate for the wider adoption of multi-camera visual tracking in robotics based on two principal advantages observed in our experiments: enhanced trajectory reliability in feature-poor environments when operating in odometry mode, and increased loop closure detection rates in SLAM mode.
+4.1 Trajectory Reliability Under Visual Occlusion
+Figure 8 demonstrates the system’s trajectory reliability through a practical robustness test.
+We equipped a Carter robot with 4 stereo cameras and subjected it to challenging visual conditions during continuous motion.
+Cameras were randomly occluded with opaque film for intervals of 20–60 seconds, with the constraint that at least one stereo pair remained unobstructed at a given moment.
+Despite these severe visual constraints, the odometry trajectory maintained stability without exhibiting jumps or discontinuities, validating the system’s robustness to partial visual obstruction.
+This capability represents a significant advancement for visual tracking systems.
+While single-camera configurations can maintain tracking in feature-poor environments only briefly—typically relying on IMU or depth sensor fusion—our multi-camera approach preserves stable and reliable visual tracking over extended periods even with degraded visual input.
 Figure 8: Occlusion resilience test results using the Carter robot.
 The figure shows the continuous trajectory successfully maintained by cuVSLAM multicamera visual odometry despite random camera occlusions using non-transparent film.
-Numbered image rows correspond to instantaneous views from the robot’s four stereo cameras (showing only the left image from each stereo pair) at the respective numbered locations along the trajectory.
-Additionally, the increased loop closure detection rate resulting from an expanded collective field of view is demonstrated in Fig.
-9, comparing the number of loop closure events between single-stereo and four-stereo camera setups for the same trajectory.
-This higher frequency of camera pose adjustments yields a quantitative improvement of approximately 40% in SLAM accuracy metrics compared to odometry mode, as demonstrated in Tab.
-2 using the Multi-Stereo R2B dataset.
+Numbered image rows correspond to instantaneous views from the robot’s 4 stereo cameras (showing only the left image from each stereo pair) at the numbered locations along the trajectory.
+4.2 Enhanced Loop Closure Detection
+The expanded collective field of view afforded by multiple cameras directly translates to increased loop closure detection rates.
+Figure 9 presents a comparative analysis of loop closure events between single-stereo and four-stereo camera configurations along identical trajectories.
+The multi-camera setup demonstrates a substantially higher frequency of successful loop closures, enabling more frequent pose graph optimization.
+This increased loop closure rate yields quantitative improvements in localization accuracy.
+As shown in Table 3, evaluation on the Multi-Stereo R2B dataset reveals approximately 40% improvement in SLAM accuracy metrics compared to pure odometry mode, underscoring the practical benefits of the multi-camera approach for robust localization.
 Figure 9: Comparison of loop closure detection events between single-camera and multi-camera (four stereo pairs) SLAM configurations on identical trajectories.
 The multi-camera setup demonstrates improved mapping consistency due to its wider cumulative field-of-view.
+5 Conclusions
+This technical report demonstrates the robust real-time performance of the cuVSLAM visual tracking library across diverse robotic platforms, including automotive vehicles, aerial drones, wheeled robots, and handheld camera systems.
+Our comprehensive evaluation encompasses both real-world deployments and synthetic datasets, validating the system’s capabilities under varied operational conditions.
+The cuVSLAM library offers several key advantages that position it as a production-ready solution for modern robotics applications:
+  • Accessibility and Integration: The library is readily available through both ISAAC ROS [^1] https://github.com/NVIDIA-ISAAC-ROS/isaac_ros_visual_slam and Python [^2] https://github.com/NVlabs/PyCuVSLAM APIs, supporting deployment on both desktop and embedded NVIDIA platforms.
+This dual-API approach ensures flexibility for developers working across different software ecosystems.
+  • Computational Efficiency: The system maintains low computational overhead while delivering high-accuracy localization, enabling seamless integration into complex robotic stacks such as NVIDIA Isaac Perceptor [^3] https://developer.nvidia.com/isaac/perceptor, which incorporates NVBLOX Millane et al.
+(2023) for 3D reconstruction and Nav2 Macenski et al.
+(2020) for autonomous navigation.
+  • Minimal Configuration Overhead: Unlike many SLAM systems that require extensive parameter tuning for different scenarios, cuVSLAM provides robust out-of-the-box performance across diverse environments and sensor configurations, significantly reducing deployment time and complexity.
+Looking forward, the demonstrated multi-camera capabilities and robust performance characteristics of cuVSLAM establish a strong foundation for next-generation autonomous systems that demand reliable visual localization in challenging real-world environments.
+The library’s proven versatility across platforms and scenarios, combined with its efficient GPU-accelerated implementation, makes it a compelling choice for both research prototypes and production robotic systems.
 Appendix A Appendix
-A.1 Appendix A
-The following publicly available datasets were used for validating the libraries:
-The TartanAir dataset includes 120 Hard sequences, the complete list is provided in Tab.
-3 .
-From the TUM RGB-D SLAM dataset, only the sequences from the freiburg3 subset were utilized, as these sequences specifically provide rectified images required by the cuVSLAM Mono-Depth mode.
-The 10 sequences employed from this subset are listed in Tab.
-4.
-Table 3: TartanAir dataset sequences used for evaluation
-Environment Sequences
-abandonedfactory P000-P011
-amusement P000-P007
-carwelding P000-P003
-endofworld P000-P002, P005-P006
-hospital P037-P039, P042, P044-P049
-japanesealley P000-P005
-neighborhood P000-P002, P004,
-P006-P010, P012-P017
-ocean P000-P008
-office2 P000-P002, P004-P008, P010
-office P000, P001, P004, P006
-oldtown P000-P008
-seasidetown P000, P002, P004
-seasonsforest P001-P002, P004-P006
-seasonsforest_winter P010-P018
-soulcity P000-P002, P004, P008, P009
-westerndesert P000, P002, P003, P005-P007
-Table 4: TUM RGB-D SLAM dataset sequences used for evaluation
-Sequences
-large cabinet validation
-long office household
-nostructure texture far
-nostructure texture near withloop
-sitting halfsphere
-sitting xyz
-sitting xyz validation
-structure texture far
-structure texture near
-teddy
-A.2 Appendix B
-The Multi-Stereo R2B dataset was collected using a wheeled Carter robot equipped with four Full HD RGB stereo Hawk cameras operating at 30 FPS.
-Stereo cameras are hardware-synchronized, with a maximum synchronization time difference of approximately 0.1 ms within each 8-image batch.
-Each recording has been validated for completeness regarding lost image batches; the maximum loss rate of image batches is below 0.04%.
-Fig.
-10 shows the trajectories of all sequences along with detailed statistics, including the trajectory distance and the number of synchronized image batches for each recording.
-Figure 10: Trajectories of sequences from the Multi-Stereo R2B dataset, along with corresponding distances and the number of 8-image batches per trajectory
+A.1 Appendix A. Detailed Description of Dataset Validation
+A.1.1 KITTI Dataset
+We evaluated our method on sequences 0-10 of the KITTI dataset, for which ground truth trajectories are available.
+For comparison with the conventional KITTI leaderboard (1), Figure 10 presents relative translation and rotation errors calculated on trajectory segments of lengths 100, 200, 300, 400, 500, 600, 700, and 800 meters.
+In contrast Table 2 presents results calculated without trajectory segmentation to maintain metric comparability across different datasets.
+Figure 10: cuVSLAM evaluation results on KITTI odometry benchmark sequences 00-10.
+Translation and rotation errors are calculated on segments of 100-800m following the KITTI public leaderboard methodology, averaged metrics: ATE: 0.85%, ARE: 0.0025 [deg/m].
+A.1.2 TUM RGB-D Dataset
+To validate the Mono-Depth tracking mode, we employed a subset of 10 sequences from the Freiburg3 collection of the TUM RGB-D dataset (20).
+These sequences were selected because they provide rectified images, which are required by the cuVSLAM Mono-Depth mode.
+We performed minimal preprocessing, implementing only depth-image pair synchronization without any additional image or depth processing.
+The complete list of evaluated sequences is presented in Table 4.
+Table 4: Results of cuVSLAM Mono-Depth Odometry and SLAM evaluation on TUM RGB-D dataset
+Sequence Mode ATE ARE APE
+large cabinet validation ODOM 1.74 2.08 0.12
+SLAM 1.53 1.60 0.11
+long office household ODOM 1.27 7.89 0.20
+SLAM 0.96 5.94 0.06
+nostructure texture far ODOM 1.29 1.66 0.07
+SLAM 1.44 1.63 0.06
+nostructure texture near withloop ODOM 1.23 4.62 0.10
+SLAM 1.25 4.71 0.04
+sitting halfsphere ODOM 2.07 3.67 0.12
+SLAM 0.61 1.42 0.05
+sitting xyz ODOM 1.03 1.21 0.05
+SLAM 0.65 0.86 0.03
+sitting xyz validation ODOM 0.90 1.42 0.05
+SLAM 0.44 0.67 0.03
+structure texture far ODOM 1.11 1.46 0.04
+SLAM 0.55 0.69 0.02
+structure texture near ODOM 1.45 2.15 0.03
+SLAM 1.45 2.09 0.03
+teddy ODOM 1.38 29.04 0.31
+SLAM 1.04 21.70 0.22
+A.1.3 TartanGround and TartanAir V2
+Odometry Mode.
+We validated cuVSLAM using 216 sequences from TartanGround  (18) and 539 sequences from TartanAir V2  (21) Hard datasets.
+Complete sequence lists are provided in Tables 5 and 9, respectively.
+All results were obtained using a four stereo-camera configuration (front, back, left, and right) with 640×640 undistorted images.
+While the TartanAir V2 dataset supports two additional cameras (top and bottom), their inclusion yielded minimal metric improvement (less than 5%).
+For the TartanGround dataset, we observed slight performance degradation when using these additional cameras, attributed to meaningless feature selection from sky regions and repetitive floor textures in planar robot navigation areas.
+To facilitate comparison with other visual tracking libraries, we calculated additional metrics—relative translation and rotation per frame—using the official TartanAir GitHub repository (2).
+As not all environments mentioned in the original TartanGround paper were available for download, we provide environment-averaged results to enable future comparative studies.
+Table 5: Odometry evaluation results for cuVSLAM Multi-Stereo configuration (4 stereo cameras) on TartanGround dataset. Values represent averages computed per environment.
+Environment Sequences ATE ARE APE $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
+Downtown P0-P17 0.30 0.41 0.11 0.0019 0.00013
+ForestEnv P0-P22 0.45 0.72 0.19 0.0024 0.00036
+Gascola P0-P21 0.24 0.46 0.09 0.0012 0.00018
+GreatMarsh P1-P19, P21-P29 2.34 4.21 0.45 0.0037 0.00126
+ModernCityDowntown P1-P10 0.19 0.39 0.08 0.0012 0.00014
+ModularNeighborhood P0-P29 0.42 0.65 0.12 0.0014 0.00016
+NordicHarbor P0-P27 0.23 0.41 0.06 0.0010 0.00015
+OldTownFall P0-P7 0.15 0.34 0.04 0.0011 0.00013
+OldTownSummer P0-P7 0.15 0.34 0.05 0.0011 0.00013
+SeasonalForestAutumn P0-P19 0.27 0.42 0.11 0.0017 0.00022
+SeasonalForestSpring P0-P19 0.23 0.49 0.12 0.0016 0.00024
+SeasonalForestWinter P0-P21 0.27 0.51 0.08 0.0011 0.00018
+Average 0.54 0.95 0.14 0.0017 0.00032
+TartanAir V2 Easy sequences were excluded from our analysis as they present minimal challenge for cuVSLAM, yielding consistently high tracking quality similar to the TartanGround dataset.
+To support this decision, Table 6 presents a comparison of averaged four-stereo-camera odometry results between Easy and Hard subsets of TartanAir V2, with example trajectory estimates shown in Figure 11.
+Table 6: Comparative odometry evaluation of cuVSLAM Multi-Stereo (4 stereo cameras) on TartanAir V2 Easy and Hard sequences. Results show average performance metrics for selected environments.
+Environment Difficulty ATE ARE APE
+AbandonedCable easy 0.05 0.56 0.26
+hard 1.93 14.5 8.65
+AbandonedFactory2 easy 0.02 0.13 0.02
+hard 0.27 1.41 0.34
+CoalMine easy 0.03 0.16 0.02
+hard 0.37 1.25 0.56
+GothicIsland easy 0.08 0.58 0.17
+hard 1.73 13.44 4.48
+Hospital easy 0.04 0.60 0.06
+hard 1.50 16.17 3.25
+JapaneseAlley easy 0.06 0.34 0.06
+hard 0.97 5.45 0.50
+OldBrickHouseDay easy 0.06 0.54 0.04
+hard 1.74 11.15 1.48
+Figure 11: Example trajectories predicted by cuVSLAM (4 Multi-Stereo Odometry) on TartanAir V2 (Easy and Hard).
+Black lines represent ground truth, green lines represent odometry predictions.
+Columns correspond to the same environment and sequence label at different difficulties.
+SLAM Mode.
+To evaluate the impact of SLAM mode on trajectory estimation, we analyzed only sequences containing loop closures.
+This resulted in a subset of 32 sequences for TartanGround and 506 sequences for TartanAir V2 Hard.
+Complete tracking results are provided in Tables 7 and 9, with all results averaged across environments.
+Table 7: Comparative evaluation of odometry and SLAM performance for cuVSLAM Multi-Stereo (4 stereo cameras) on the TartanGround dataset.
+To assess the impact of SLAM mode, a subset of sequences containing looped trajectories with observable loop closures was selected.
+Reported values are averages computed per environment.
+Environment Sequences Mode ATE ARE APE
+Downtown P7 ODOM 0.35 0.59 0.09
+SLAM 0.25 0.41 0.09
+ForestEnv P1, P14, P17, P21 ODOM 0.42 0.78 0.16
+SLAM 0.38 0.61 0.17
+Gascola P0, P3-P6, P9 ODOM 0.17 0.41 0.08
+SLAM 0.13 0.31 0.06
+ModernCityDowntown P2, P5-P7 ODOM 0.17 0.32 0.08
+SLAM 0.12 0.24 0.06
+ModularNeighborhood P12, P20 ODOM 0.42 1.17 0.14
+SLAM 0.35 1.02 0.05
+NordicHarbor P1, P11, P17-P19 P23, P27 ODOM 0.17 0.41 0.08
+SLAM 0.10 0.26 0.04
+OldTownFall P2 ODOM 0.07 0.31 0.05
+SLAM 0.12 0.26 0.06
+OldTownSummer P2-P4 ODOM 0.09 0.29 0.06
+SLAM 0.06 0.25 0.03
+SeasonalForestAutumn P5, P8, P17 ODOM 0.20 0.32 0.07
+SLAM 0.19 0.33 0.07
+SeasonalForestSpring P9 ODOM 0.15 0.49 0.05
+SLAM 0.11 0.39 0.02
+Average ODOM 0.21 0.48 0.088
+SLAM 0.17 0.37 0.065
+For the TartanAir V2 dataset, most environments utilized all available sequences.
+Sequences excluded from SLAM validation are marked with an asterisk (P*), while their corresponding visual odometry results are labeled ODOM*, indicating performance metrics computed across all available sequences for those specific environments, including the excluded sequences.
+A.1.4 Mutli-Stereo R2B Dataset
+The Multi-Stereo R2B dataset was acquired using a wheeled Carter robot platform equipped with four stereo Hawk cameras, each capturing RGB images at a resolution of 1920×1200 pixels and 30 frames per second.
+The stereo camera pairs are hardware-synchronized to ensure temporal coherence, with synchronization precision maintained within 0.1 ms for each 8-image batch.
+Data integrity validation confirmed minimal frame loss across all recordings, with the maximum batch loss rate remaining below 0.04%.
+Figure 12 illustrates the complete set of recorded trajectories, along with corresponding trajectory lengths and synchronized image batch counts for each sequence.
+Figure 12: Trajectories predicted by cuVSLAM (4 Multi-Stereo Odometry) on the proprietary R2B wheeled robot dataset.
+Black lines represent ground truth; green lines represent odometry predictions.
+Attached table lists trajectory distances, number of 8-image batches, and trajectory numbering per environment (left to right).
+A.2 Appendix B: Hardware Scalability on Embedded Platforms
+This appendix presents an analysis of hardware utilization and visual tracking performance for cuVSLAM on embedded platforms, comparing Multi-Stereo and Mono-Depth modes as representative examples of sparse and dense feature-based approaches, respectively.
+Experimental Setup.
+The primary hardware configuration consisted of an NVIDIA Jetson Orin AGX operating in MAXN mode with 3 Intel RealSense D435 stereo cameras connected via USB and hardware-synchronized.
+Different resolutions were obtained directly from the cameras by modifying their configuration during the initialization step of each experiment.
+For Full HD resolution testing, we employed a set of 3 RGB Hawk stereo cameras connected via FAKRA2 connectors with hardware synchronization.
+All experiments were conducted in real-time with live camera feeds connected directly to the Jetson boards, and cuVSLAM calculations were performed on the same board in real-time.
+The system operated in Performance mode, utilizing the Isaac ROS (ROS2) stack for image acquisition and Multi-Stereo visual tracking, while the Python stack was employed for Mono-Depth visual tracking and processing.
+Methodology.
+Resource utilization was measured through a two-phase experimental protocol designed to isolate the computational overhead of cuVSLAM.
+In the first phase, system utilization was logged while receiving frames from freely moving cameras without invoking visual tracking, establishing a baseline measurement.
+In the second phase, the identical setup was used with cuVSLAM visual tracking enabled.
+CPU and GPU utilization metrics were sampled at 15ms intervals, and the time spent on cuVSLAM frame processing was logged for each frame and subsequently averaged.
+The net computational resource utilization of cuVSLAM was calculated by subtracting the baseline measurements from the active tracking measurements.
+Jetson Orin AGX Performance.
+Figure 13 illustrates the significant computational efficiency advantage of the sparse feature-based approach for tracking, demonstrating 3-7× faster processing times for 6-image batches compared to single image-depth pairs.
+The phenomenon where the 3-stereo camera visual tracking in Full HD configuration exhibited lower hardware utilization than HD configuration is attributed to differences in the camera field-of-view, with Hawk cameras provided 121.5° × 73.5° FOV in FHD mode versus RealSense D435 cameras provided 87° × 58° FOV in HD mode.
+The wider field-of-view affects the lifetime of keypoints as described in Section 2.1, triggering more frequent keyframe creation and consequently invoking computationally intensive operations including 3D landmark triangulation and sparse bundle adjustment, as detailed in Sections 2.1 and  2.2.
+Figure 13: cuVSLAM realtime network hardware utilization and callback time per input image on Jetson AGX Orin as a function of resolution for 3-Stereo Camera Visual Odometry (left) and Mono-Depth Visual Odometry (right) at 30 FPS.
+For further details, refer to Appendix  A.2
+For Mono-Depth mode, the average callback time for 720p resolution reached 37.7ms, which exceeds the 33.3ms frame interval required for 30 FPS operation.
+This timing constraint indicates GPU overload on the Orin AGX platform in this configuration, and according to our measurements, resulted in the system skipping approximately half of the incoming frames.
+Consequently, for HD resolution applications, we recommend either reducing the frame rate to 15 FPS or decreasing the resolution to VGA, as our experiments revealed no significant difference in cuVSLAM visual tracking quality between VGA and HD resolutions.
+Jetson Orin Nano Super Performance.
+To evaluate the feasibility of deploying cuVSLAM on entry-level hardware, we conducted experiments replacing the Jetson Orin AGX with a Jetson Orin Nano Super operating in 25W mode and utilizing exclusively the Python stack.
+The experiments explore possibility of using multi-stereo odometry mode for 640x360 resolution and 30 FPS followed the methodology described above with free camera movement.
+Figure 14: Hardware utilization and callback time on Jetson Orin Nano for Multi-Stereo Visual Odometry using RealSense cameras at 640×360 resolution and 30 FPS.
+Results are presented for varying numbers of cameras, showing total utilization (faded color bars) and net cuVSLAM utilization (normal color bars).
+The results of average net cuVSLAM resource utilization, total system load, and average processing time are presented in Figure 14.
+The findings demonstrate substantial resource headroom, enabling the use of higher frame rates for configurations with one or 2 stereo cameras.
+For Mono-Depth processing at 30 FPS, stable operation was achieved only at the minimum resolution of 424×240 pixels.
+Although the average GPU load remained below 40%, the observed average cuVSLAM call time of 31ms approached the critical 33ms frame interval at 30 FPS.
+When combined with additional operations in the image acquisition loop, this resulted in exceeding the frame interval and consequently dropping approximately 30% of incoming frames.
+It is important to note that for scenarios involving ground robots, the observed computational load will be lower than our free-movement test scenarios due to the reduced number of keyframes generated during predominantly inertial movements.
+However, for systems involving freely moving cameras with rapid scene changes, such as handheld cameras or unmanned aerial vehicles, we recommend increasing frame rates to 60 FPS when using RealSense cameras to ensure robust tracking performance.
+The selection of appropriate resolution and frame rate parameters should be carefully balanced based on the available computational resources and specific application requirements.
+Table 8: TartanAir V2 cuVSLAM evaluation results (Envs 1–35, see Appendix A.1.3 for details)
+Env Seq Mode ATE ARE APE
+1 Abandoned- Cable P0-P7 ODOM 1.90 13.77 8.97
+SLAM 1.54 11.47 7.15
+2 Abandoned- Factory P1-P6 P8-P9 ODOM 1.07 4.89 1.06
+SLAM 0.77 3.70 0.95
+3 Abandoned- Factory2 P0-P5 ODOM 0.27 1.41 0.34
+SLAM 0.19 0.84 0.24
+4 Abandoned- School P0-P9 ODOM 3.59 33.78 5.41
+SLAM 3.31 31.42 5.12
+5 American- Diner P0, P2* P3-P4 ODOM* 5.53 13.27 1.25
+ODOM 5.21 11.12 1.05
+SLAM 4.98 10.94 0.93
+6 Amusement- Park P0, P1* P2-P3, P4* P5-P7 ODOM* 1.15 3.74 1.53
+ODOM 1.36 5.15 1.71
+SLAM 1.36 5.49 1.71
+7 Ancient- Towns P0-P6 ODOM 2.86 10.98 2.52
+SLAM 2.50 8.93 2.23
+8 Antiquity3D P1-P5, P6* P7, P8*,P9 ODOM* 2.30 17.90 4.05
+ODOM 1.95 16.53 3.69
+SLAM 1.84 15.23 3.55
+9 Apocalyptic P0-P4 ODOM 0.92 4.25 1.25
+SLAM 0.66 2.81 0.98
+10 ArchVizTiny- HouseDay P0, P1* P2-P3,P4* P5-P6 ODOM* 3.75 4.02 0.37
+ODOM 2.69 3.48 0.27
+SLAM 2.44 3.14 0.25
+11 ArchVizTiny- HouseNight P0*, P1 P5, P6* P2-P4* ODOM* 5.37 6.23 0.47
+ODOM 4.19 5.64 0.39
+SLAM 4.15 5.91 0.36
+12 Brushify- Moon P1-P5 ODOM 5.64 40.15 44.55
+SLAM 6.23 41.28 45.50
+13 CarWelding P0-P8 ODOM 1.14 5.74 1.34
+SLAM 0.94 4.51 1.27
+14 Castle- Fortress P0-P11 ODOM 0.93 7.33 2.65
+SLAM 0.90 6.41 2.50
+15 CoalMine P0-P5 ODOM 0.34 1.42 0.51
+SLAM 0.20 0.67 0.38
+16 Construct- Site P1, P3* P5* P6-P11 ODOM* 2.80 8.81 3.15
+ODOM 2.90 9.23 3.31
+SLAM 2.80 8.74 3.14
+17 Country- House P0-P5 ODOM 0.87 2.69 0.20
+SLAM 0.16 0.62 0.05
+18 CyberPunk- Downtown P0-P6 ODOM 1.19 7.27 4.01
+SLAM 1.27 8.22 3.98
+19 Cyberpunk P0-P7 ODOM 1.13 4.65 3.38
+SLAM 1.13 5.43 4.23
+20 Desert- GasStation P0-P4 ODOM 2.44 7.11 2.53
+SLAM 2.31 6.78 2.43
+21 Downtown P0-P8 ODOM 1.20 5.66 3.59
+SLAM 1.21 5.52 3.35
+22 Factory- Weather P1-P7, P8* P9-P12 ODOM* 6.23 39.06 22.37
+ODOM 6.31 40.09 22.64
+SLAM 6.17 37.40 22.22
+23 Fantasy P0-P6 ODOM 1.01 4.16 1.45
+SLAM 1.01 4.70 1.87
+24 ForestEnv P0, P2-P5 P7, P9 P10*, P11 ODOM* 3.67 21.93 13.76
+ODOM 3.55 21.75 14.42
+SLAM 3.62 23.64 15.47
+25 Gascola P0-P9 ODOM 2.36 14.83 5.47
+SLAM 2.49 14.56 6.13
+26 Gothic- Island P0-P10 ODOM 1.68 13.19 4.00
+SLAM 1.44 11.60 3.54
+27 HQWestern- Saloon P0-P7 ODOM 2.67 16.93 2.01
+SLAM 1.96 8.81 1.60
+28 HongKong P0, P1* P2-P4 ODOM* 5.26 18.85 5.02
+ODOM 4.81 18.85 3.99
+SLAM 4.75 17.78 3.89
+29 Hospital P0-P10 ODOM 1.50 16.17 3.25
+SLAM 1.30 14.85 2.93
+30 House P0-P7 ODOM 1.47 9.80 0.47
+SLAM 0.98 6.43 0.26
+31 Industrial- Hangar P0-P6 ODOM 3.43 19.61 8.21
+SLAM 3.48 19.86 8.03
+32 Japanese- Alley P0-P6 ODOM 0.94 5.27 0.48
+SLAM 0.34 1.87 0.39
+33 Japanese- City P0-P10 ODOM 1.81 7.89 2.08
+SLAM 1.74 7.80 2.03
+34 MiddleEast P0-P10 ODOM 2.96 13.18 2.64
+SLAM 2.65 11.34 2.47
+35 Mod- UrbanCity P0-P11 ODOM 3.65 23.51 5.74
+SLAM 3.70 23.22 5.60
+Table 9: TartanAir V2 cuVSLAM evaluation results (Envs 36–71, see Appendix A.1.3 for details)
+Env Seq Mode ATE ARE APE
+36 ModernCity Downtown P0-P8 ODOM 2.34 11.70 4.01
+SLAM 2.08 11.34 3.60
+37 Modular- Neighborhood P0-P10 ODOM 4.14 22.48 17.10
+SLAM 3.60 21.04 13.85
+38 ModularNeigh borhoodIntExt P1-P5 P7-P8 ODOM 2.64 15.80 4.74
+SLAM 2.62 13.93 5.13
+39 NordicHarbor P0-P7 ODOM 2.07 14.37 6.04
+SLAM 1.97 15.04 5.77
+40 Ocean P1-P3 P4* P5-P7 ODOM* 1.76 8.13 1.99
+ODOM 1.65 8.66 2.36
+SLAM 1.47 7.12 2.22
+41 Office P0-P6 ODOM* 3.51 16.02 2.71
+SLAM 3.36 14.84 2.49
+42 OldBrick- HouseDay P1-P2 P3* P4-P7 ODOM* 1.74 11.15 1.48
+ODOM 1.54 11.80 1.25
+SLAM 1.29 7.76 1.10
+43 OldBrick- HouseNight P1*, P2 P4-P7 ODOM* 1.83 16.07 1.53
+ODOM 1.68 16.35 1.56
+SLAM 1.59 13.05 1.50
+44 Old- IndustrialCity P0-P9 ODOM 2.25 8.60 4.33
+SLAM 1.58 7.20 3.76
+45 Old- Scandinavia P1-P6 ODOM* 2.22 11.68 8.90
+SLAM 2.15 11.47 8.38
+46 OldTownFall P0*,P1* P2 ODOM* 1.62 6.44 1.53
+ODOM 1.25 4.09 0.11
+SLAM 0.75 2.47 0.28
+47 OldTownNight P0-P2 ODOM 0.83 4.74 1.24
+SLAM 0.61 3.49 0.95
+48 OldTown- Summer P0-P2 ODOM 2.42 11.90 1.57
+SLAM 2.41 11.84 1.54
+49 OldTown- Winter P0-P2 ODOM 0.60 2.14 0.72
+SLAM 0.53 2.33 0.75
+50 PolarSciFi P0-P8 ODOM 2.76 12.74 2.19
+SLAM 2.68 12.34 2.18
+51 Prison P0-P11 ODOM 4.14 21.37 4.59
+SLAM 4.12 20.78 4.68
+52 Restaurant P1-P8 ODOM 1.51 10.99 0.97
+SLAM 1.03 8.00 0.70
+53 RetroOffice P0-P5 ODOM 1.14 3.11 0.10
+SLAM 0.94 2.85 0.09
+54 Ruins P1-P6 P8-P9 ODOM 3.07 16.68 9.43
+SLAM 3.05 17.07 9.09
+55 SeasideTown P0-P7 ODOM 0.67 1.79 0.56
+SLAM 0.61 1.58 0.44
+56 SeasonForest- Autumn P0-P2 ODOM 5.41 35.20 21.38
+SLAM 5.10 31.76 18.73
+57 SeasonForest- Spring P0-P3 ODOM 4.09 25.14 15.03
+SLAM 3.93 28.30 15.02
+58 SeasonForest- SummerNight P0-P1 ODOM 3.65 21.45 25.05
+SLAM 3.45 19.42 23.00
+59 SeasonForest- Winter P0-P2 ODOM 6.73 30.59 15.47
+SLAM 7.53 32.61 19.11
+60 SeasonForest- WinterNight P0-P2 ODOM 4.77 30.45 27.40
+SLAM 5.14 30.09 26.31
+61 Sewerage P1-P11 ODOM 3.94 30.11 6.74
+SLAM 3.87 27.64 5.92
+62 ShoreCaves P1-P8 ODOM 1.86 7.56 3.41
+SLAM 1.77 7.20 3.00
+63 Slaughter P0, P1* P2-P10 P11* ODOM* 4.52 26.89 6.04
+ODOM 3.97 27.19 5.53
+SLAM 3.80 24.49 5.34
+64 SoulCity P0*,P1,P2* P3-P4, P5* P6, P7 ODOM* 5.22 32.59 7.30
+ODOM 4.74 28.79 7.96
+SLAM 4.82 29.27 8.10
+65 Supermarket P0*,P1,P2 P3*, P4-P6 ODOM* 4.70 23.19 3.69
+ODOM 4.82 22.64 3.66
+SLAM 4.73 23.38 3.60
+66 Terrain- Blending P1, P3-P4 ODOM 0.45 1.65 1.31
+SLAM 0.55 1.90 1.55
+67 Urban- Construction P0-P4 ODOM 0.35 1.76 1.30
+SLAM 0.31 1.50 1.24
+68 VictorianStreet P0-P7 ODOM 0.99 4.67 0.43
+SLAM 0.72 3.17 0.25
+69 WaterMillDay P1-P3, P6 ODOM 0.48 2.54 0.74
+SLAM 0.30 1.12 0.59
+70 WaterMillNight P0-P6 ODOM 0.34 0.94 0.78
+SLAM 0.29 0.92 0.72
+71 Western- DesertTown P1-P7, P9 P11, P13 ODOM 1.89 13.77 8.09
+SLAM 1.78 12.97 7.32

### v3

@@
 Besides the map, the frontend consists of the 2D module, responsible for feature selection, 2D feature tracking, and keyframe selection, and the 3D module, which performs pose estimation based on the local map, 2D information, and other data sources (e.g., IMU or depth measurements).
-In contrast to the frontend, cuVSLAM’s backend focuses on global map and pose consistency.
+In contrast to the frontend, cuVSLAM backend focuses on global map and pose consistency.
 It processes frontend outputs and builds a globally consistent map that integrates camera poses, 2D observations, 3D landmarks, pose deltas, and visual features.
@@
 Edge direction describes the cross-camera tracking between adjacent cameras, i.e., if $e_{ij}$ exists, cuVSLAM performs 2D tracking from camera $c_{i}$ to $c_{j}$.
-The FIG is built automatically at cuVSLAM’s startup.
+The FIG is built automatically at cuVSLAM startup.
 Provided with a list of camera extrinsics/intrinsics, cuVSLAM iterates over each pair of cameras and creates a new edge in the graph when cameras have overlapping fields of view (FoV).
@@
 Figure 6: Factor graphs for VIO pipeline.
-cuVSLAM’s loop closing and global map refinement module helps limit global drift by enforcing global consistency of the map.
+cuVSLAM loop closing and global map refinement module helps limit global drift by enforcing global consistency of the map.
 It consists of 6 DoF keyframe poses, landmarks and their observations, pose deltas, and sparse image features.
@@
 3.1 System Performance and Hardware Utilization
-We evaluated cuVSLAM’s performance through two complementary approaches: real-time execution on embedded hardware to measure resource utilization and single-frame processing time, and offline processing of warehouse environment recordings on workstation hardware.
+We evaluated cuVSLAM performance through two complementary approaches: real-time execution on embedded hardware to measure resource utilization and single-frame processing time, and offline processing of warehouse environment recordings on workstation hardware.
 The test scenarios involved processing feature-rich images captured from either a wheeled robot navigating a warehouse environment or handheld camera movements.
@@
 Table 1 summarizes the key performance metrics, with comprehensive results provided in Appendix A.2.
-The left portion of Table 1 presents execution times for cuVSLAM’s visual tracking function, excluding image acquisition and post-processing overhead.
+The left portion of Table 1 presents execution times for cuVSLAM visual tracking function, excluding image acquisition and post-processing overhead.
 These measurements were obtained using recorded color images at 768×480 resolution (with the exception of mono-depth mode at 640×480 resolution) on two distinct hardware configurations: a desktop workstation featuring an Intel i7-14700 CPU with NVIDIA RTX 4090 GPU, and an embedded Jetson Orin AGX 64GB single-board computer operating in MAXN power mode.
@@
 The reduced framerate was necessary to ensure valid IMU-based integration, providing 6-7 IMU measurements between each stereo image pair.
-To accurately isolate cuVSLAM’s hardware requirements, we employed a controlled experimental methodology: we first established baseline system utilization with the RealSense camera in motion but without cuVSLAM processing, then repeated the similar camera movement pattern with cuVSLAM visual tracking active.
+To accurately isolate cuVSLAM hardware requirements, we employed a controlled experimental methodology: we first established baseline system utilization with the RealSense camera in motion but without cuVSLAM processing, then repeated the similar camera movement pattern with cuVSLAM visual tracking active.
 Resource utilization metrics were sampled at 15 ms intervals for equal durations in both experiments.
@@
 This methodology provides an accurate assessment of the computational overhead introduced by the library.
-Additional evaluation details, including performance benchmarks for cuVSLAM executed via Python API on entry-level Jetson Orin Nano hardware, are presented in Appendix A.2.
+Additional details on evaluation and an exploration of hardware utilization across different resolutions and modes on Jetson devices are provided in Appendix A.2.
 3.2 Benchmarking
-We evaluated cuVSLAM’s different operation modes using several publicly available datasets.
-The Mono-Depth mode was validated on the ICL-NUIM Handa et al.
+We evaluated cuVSLAM different operation modes using several publicly available datasets.
+The Mono-Depth mode was validated on the AR-table Chen et al.
+(2023), ICL-NUIM Handa et al.
 (2014) and TUM RGB-D Sturm et al.
@@
 (2013) datasets.
-The Stereo-Inertial mode was assessed using the EuRoC dataset.
+The Stereo-Inertial mode was assessed using the EuRoC Burri et al.
+(2016) and TUM-VI Schubert et al.
+(2018) datasets.
 The Multi-Stereo mode was evaluated on simulated TartanAir V2 Wang et al.
@@
 (2025) datasets and our proprietary R2B real-world dataset.
-To quantify trajectory estimation accuracy, we employed standard metrics: Average Relative Translation and Rotation Errors Geiger et al.
-(2013), and Absolute Pose Error with trajectory alignment and scale correction Sturm et al.
-(2012) using the EVO library Grupp (2017).
 Table 2: Evaluation results for cuVSLAM Visual Odometry and SLAM across various operating modes
-Mode Dataset Method ATE, % ARE, deg APE (RMSE)
-Mono-Depth ICL-Nuim ORBSLAM3 - - -
+Mode Dataset Method avgRTE, % avgRE, deg RMSE APE
+Mono-Depth AR table ORBSLAM3 - - -
+DPVO 4.42 1.65 0.77
+cuVSLAM Odom 0.34 3.59 0.09
+cuVSLAM SLAM 0.19 1.68 0.025
+ICL-Nuim ORBSLAM3 - - -
 DPVO 10.28 0.77 0.61
@@
 cuVSLAM SLAM 0.29 2.27 0.13
+TUM-VI Room ORBSLAM3 2.17 1.37 0.077
+DPVO - - -
+cuVSLAM Odom 0.20 3.85 0.18
+cuVSLAM SLAM 0.12 3.00 0.12
+To quantify visual tracking accuracy, we employed standard metrics: Average Relative Translation (avgRTE) and Rotation (avgRE) Errors Geiger et al.
+(2013), and Absolute Pose Error (RMSE APE) with trajectory alignment and scale correction Sturm et al.
+(2012) using the EVO library Grupp (2017).
 Table 2 presents comprehensive benchmarking results across popular datasets from various domains.
@@
 Since scale corrections were applied to final trajectories rather than successive pose estimations, translation results may not be meaningful for some datasets; therefore, relative rotation error should be considered the primary metric.
+Figure 7: Sample camera views from the R2B dataset captured across diverse environments including large office spaces, industrial warehouses, and narrow corridors in both office and storage areas.
+The images show simultaneous captures from 4 stereo cameras (displaying only the left image from each stereo pair).
 To ensure accurate and representative comparisons, we excluded clearly identified outlier and corrupted sequences.
@@
 Additionally, many sequences lack overlapping trajectories, preventing loop closure detection—a critical component for evaluating SLAM performance versus odometry.
-Consequently, we validated SLAM mode on selected subsets of TartanGround (see Table 7) and TartanAir V2 (see Tables 9 and 9) datasets.
-Figure 7: Sample camera views from the R2B dataset captured across diverse environments including large office spaces, industrial warehouses, and narrow corridors in both office and storage areas.
-The images show simultaneous captures from 4 stereo cameras (displaying only the left image from each stereo pair).
-While the new Tartan datasets provide free camera and quadruped robot movements in diverse simulated environments, we needed to rigorously validate cuVSLAM’s multi-camera capabilities in real-world scenarios.
-We therefore collected a proprietary Multi-Stereo R2B dataset in office and warehouse environments using the Nova Carter wheeled robot equipped with four stereo cameras.
+Consequently, we validated SLAM mode on selected subsets of TartanGround (see Table 9) and TartanAir V2 (see Tables 11 and 11) datasets.
+While the new Tartan datasets provide free camera and quadruped robot movements in diverse simulated environments, rigorous validation of cuVSLAM multi-camera capabilities in real-world scenarios was required.
+Therefore, a proprietary Multi-Stereo R2B dataset was collected in office and warehouse environments using the Nova Carter wheeled robot equipped with four stereo cameras.
 The example views from the R2B dataset are shown in Fig.
@@
 (2024) or employ software synchronization that cannot ensure simultaneous image capture Yin et al.
-(2021); Zhang et al.
-(2023).
-Table 3 presents cuVSLAM’s results in Multi-Stereo Camera Mode.
-Notably, results achieved on the real R2B dataset closely match those from similar domains in the TartanGround dataset, confirming the validity of both datasets and results.
-For TartanAir V2, we performed detailed validation only on the Hard dataset, which presents greater challenges due to fast six-degrees-of-freedom motion and higher translational and angular velocities.
-Results for the Easy mode are comparable to those for TartanGround, with a detailed comparison available in Table 6.
-While TartanAir V2 results indicate room for improvement, they also demonstrate the importance of multi-camera mode for robust tracking in most environments, with some showing 2–4× improvement in multi-camera SLAM mode.
-Table 3: Evaluation results for cuVSLAM Multi-Stereo Visual Odometry and SLAM
-Dataset Method ATE, % ARE, deg APE (RMSE)
+(2021); Zhang (2023).
+Table 3: Evaluation results for cuVSLAM Multi-Stereo Visual Odometry and SLAM on multi-camera datasets
+Dataset Method avgRTE, % avgRE, deg RMSE APE
 R2B cuVSLAM Odom 0.18 1.15 0.28
@@
 cuVSLAM SLAM 2.26 12.76 4.99
+Table 3 presents cuVSLAM results in Multi-Stereo Mode.
+Notably, results achieved on the real R2B dataset closely match those from similar domains in the TartanGround dataset, confirming the validity of both datasets and results.
+For TartanAir V2, we performed detailed validation only on the Hard dataset, which presents greater challenges due to fast six-degrees-of-freedom motion and higher translational and angular velocities.
+Results for the Easy mode are comparable to those for TartanGround, with a detailed comparison available in Table 8.
+While TartanAir V2 results indicate room for improvement, they also demonstrate the importance of multi-camera mode for robust tracking in most environments, with some showing 2–4× improvement in multi-camera SLAM mode.
 We anticipate that the recent growth in availability and adoption of synchronized multi-stereo camera systems from various vendors will encourage the creation of new public multi-camera datasets, facilitating further development, testing, and implementation of multi-camera SLAM systems in robotic applications.
@@
 Cameras were randomly occluded with opaque film for intervals of 20–60 seconds, with the constraint that at least one stereo pair remained unobstructed at a given moment.
+Figure 8: Occlusion resilience test using the Carter robot.
+cuVSLAM maintains continuous trajectory despite random camera occlusions.
+Numbered rows show instantaneous views from the robot’s 4 stereo cameras (left images only) at corresponding trajectory locations
 Despite these severe visual constraints, the odometry trajectory maintained stability without exhibiting jumps or discontinuities, validating the system’s robustness to partial visual obstruction.
@@
 While single-camera configurations can maintain tracking in feature-poor environments only briefly—typically relying on IMU or depth sensor fusion—our multi-camera approach preserves stable and reliable visual tracking over extended periods even with degraded visual input.
-Figure 8: Occlusion resilience test results using the Carter robot.
-The figure shows the continuous trajectory successfully maintained by cuVSLAM multicamera visual odometry despite random camera occlusions using non-transparent film.
-Numbered image rows correspond to instantaneous views from the robot’s 4 stereo cameras (showing only the left image from each stereo pair) at the numbered locations along the trajectory.
 4.2 Enhanced Loop Closure Detection
@@
 We evaluated our method on sequences 0-10 of the KITTI dataset, for which ground truth trajectories are available.
-For comparison with the conventional KITTI leaderboard (1), Figure 10 presents relative translation and rotation errors calculated on trajectory segments of lengths 100, 200, 300, 400, 500, 600, 700, and 800 meters.
+For comparison with the conventional KITTI leaderboard (1), Figure 10 presents relative translation (avgRTE) and rotation (avgRE) errors calculated on trajectory segments of lengths 100, 200, 300, 400, 500, 600, 700, and 800 meters.
 In contrast Table 2 presents results calculated without trajectory segmentation to maintain metric comparability across different datasets.
 Figure 10: cuVSLAM evaluation results on KITTI odometry benchmark sequences 00-10.
-Translation and rotation errors are calculated on segments of 100-800m following the KITTI public leaderboard methodology, averaged metrics: ATE: 0.85%, ARE: 0.0025 [deg/m].
-A.1.2 TUM RGB-D Dataset
-To validate the Mono-Depth tracking mode, we employed a subset of 10 sequences from the Freiburg3 collection of the TUM RGB-D dataset (20).
-These sequences were selected because they provide rectified images, which are required by the cuVSLAM Mono-Depth mode.
-We performed minimal preprocessing, implementing only depth-image pair synchronization without any additional image or depth processing.
-The complete list of evaluated sequences is presented in Table 4.
+Translation and rotation errors are calculated on segments of 100-800m following the KITTI public leaderboard methodology, averaged metrics: Translation: 0.85%, Rotation: 0.0025 [deg/m].
+A.1.2 TUM-VI Room Dataset
+To validate the stereo-inertial tracking mode on real hardware under aggressive camera motions, we evaluated cuVSLAM on the Room sequences of the TUM-VI dataset (20).
+These sequences were selected because they provide complete ground truth trajectories.
+We used undistorted images with a resolution of 512x512 pixels and applied a 50-pixel mask along each frame edge to mitigate significant fisheye lens distortion, which negatively affects tracking performance.
+To facilitate comparison with other visual tracking methods, we followed the evaluation protocol described in the original TUM-VI paper (20), computing the average RMSE of the Relative Pose Error (RPE) in meters/degrees over 1-second segments.
+Detailed results are provided in Table 4.
+Table 4: cuVSLAM Stereo Inertial tracking validation results on TUM-VI Room dataset RMSE RPE (m / deg) on 1 second segments
+Sequence ODOM SLAM
+room1 0.023 / 0.89 0.025 / 0.90
+room2 0.048 / 2.28 0.048 / 2.29
+room3 0.047 / 1.93 0.048 / 1.95
+room4 0.015 / 0.54 0.015 / 0.55
+room5 0.019 / 0.56 0.021 / 0.57
+room6 0.021 / 0.57 0.021 / 0.58
+A.1.3 AR-table Dataset
+This dataset contains sequences captured using an Intel RealSense D455 depth camera, providing image and depth data at a resolution of 848x480 pixels.
+The enhanced image and depth quality allows achieving performance on this real-world camera close to that obtained on the ideal simulated ICL-NUIM dataset, as shown in Table 2.
+However, consistent gaps in depth values were observed along the sides and bottom edges of the depth images.
+To mitigate this issue, we applied static masking of 24 pixels on each side and 10 pixels at the bottom.
+Additionally, several prolonged depth-image dropouts (lasting more than one second) occurred in sequences 3, 4, 5, and 8.
+To address these interruptions, we segmented the affected sequences into continuous portions, executed tracking independently on each segment, and subsequently merged the results.
+To facilitate comparison with other visual tracking methods, we followed the evaluation methodology described in the original AR-table dataset paper (8).
+Specifically, we computed Absolute Trajectory Error (ATE), reporting translation errors in centimeters and rotation errors in degrees.
+The obtained results are summarized in Table 5.
+Table 5: cuVSLAM Mono-Depth Odometry and SLAM validation results on AR table dataset ATE (degree / cm)
+Mode table_01 table_02 table_03 table_04 table_05 table_06 table_07 table_08
+ODOM 3.77 / 2.03 9.36 / 5.41 6.09 / 2.28 3.13 / 2.21 2.27 / 1.97 3.43 / 3.39 18.5 / 6.17 13.2 / 8.18
+SLAM 2.98 / 1.76 2.23 / 1.77 2.11 / 0.87 2.12 / 1.64 1.24 / 1.22 1.93 / 1.23 8.88 / 3.82 4.83 / 3.69
+A.1.4 TUM RGB-D Dataset
+To validate the Mono-Depth mode on this dataset, a subset of 10 sequences from the Freiburg3 collection was used.
+These sequences were selected as they provide rectified images, which are required by cuVSLAM Mono-Depth mode.
+Minimal preprocessing was performed, consisting only of depth-image pair synchronization without further image or depth processing.
+The complete list of evaluated sequences is provided in Table 6.
 Table 6: Results of cuVSLAM Mono-Depth Odometry and SLAM evaluation on TUM RGB-D dataset
-Sequence Mode ATE ARE APE
+Sequence Mode avgRTE avgRE RMSE APE
 large cabinet validation ODOM 1.74 2.08 0.12
@@
 SLAM 1.04 21.70 0.22
-A.1.3 TartanGround and TartanAir V2
+A.1.5 TartanGround and TartanAir V2
 Odometry Mode.
 We validated cuVSLAM using 216 sequences from TartanGround  (19) and 539 sequences from TartanAir V2  (23) Hard datasets.
-Complete sequence lists are provided in Tables 5 and 9, respectively.
+Complete sequence lists are provided in Tables 7 and 11, respectively.
 All results were obtained using a four stereo-camera configuration (front, back, left, and right) with 640×640 undistorted images.
@@
 For the TartanGround dataset, we observed slight performance degradation when using these additional cameras, attributed to meaningless feature selection from sky regions and repetitive floor textures in planar robot navigation areas.
-To facilitate comparison with other visual tracking libraries, we calculated additional metrics—relative translation and rotation per frame—using the official TartanAir GitHub repository (2).
-As not all environments mentioned in the original TartanGround paper were available for download, we provide environment-averaged results to enable future comparative studies.
+To facilitate comparison with other visual tracking libraries, we computed additional metrics—relative translation ($t_{\mathrm{rel}}$) and rotation ($r_{\mathrm{rel}}$) per frame—using scripts from the official TartanAir GitHub repository (2).
+Since not all environments from the original TartanGround paper were available, we present environment-averaged results in Table 7 to enable future comparative studies.
 Table 7: Odometry evaluation results for cuVSLAM Multi-Stereo configuration (4 stereo cameras) on TartanGround dataset. Values represent averages computed per environment.
-Environment Sequences ATE ARE APE $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
+Environment Sequences avgRTE avgRE RMSE APE $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
 Downtown P0-P17 0.30 0.41 0.11 0.0019 0.00013
@@
 Table 8: Comparative odometry evaluation of cuVSLAM Multi-Stereo (4 stereo cameras) on TartanAir V2 Easy and Hard sequences. Results show average performance metrics for selected environments.
-Environment Difficulty ATE ARE APE
+Environment Difficulty avgRTE avgRE RMSE APE
 AbandonedCable easy 0.05 0.56 0.26
@@
 This resulted in a subset of 32 sequences for TartanGround and 506 sequences for TartanAir V2 Hard.
-Complete tracking results are provided in Tables 7 and 9, with all results averaged across environments.
-Table 7: Comparative evaluation of odometry and SLAM performance for cuVSLAM Multi-Stereo (4 stereo cameras) on the TartanGround dataset.
-To assess the impact of SLAM mode, a subset of sequences containing looped trajectories with observable loop closures was selected.
-Reported values are averages computed per environment.
-Environment Sequences Mode ATE ARE APE
-Downtown P7 ODOM 0.35 0.59 0.09
-SLAM 0.25 0.41 0.09
-ForestEnv P1, P14, P17, P21 ODOM 0.42 0.78 0.16
-SLAM 0.38 0.61 0.17
-Gascola P0, P3-P6, P9 ODOM 0.17 0.41 0.08
-SLAM 0.13 0.31 0.06
-ModernCityDowntown P2, P5-P7 ODOM 0.17 0.32 0.08
-SLAM 0.12 0.24 0.06
-ModularNeighborhood P12, P20 ODOM 0.42 1.17 0.14
-SLAM 0.35 1.02 0.05
-NordicHarbor P1, P11, P17-P19 P23, P27 ODOM 0.17 0.41 0.08
-SLAM 0.10 0.26 0.04
-OldTownFall P2 ODOM 0.07 0.31 0.05
-SLAM 0.12 0.26 0.06
-OldTownSummer P2-P4 ODOM 0.09 0.29 0.06
-SLAM 0.06 0.25 0.03
-SeasonalForestAutumn P5, P8, P17 ODOM 0.20 0.32 0.07
-SLAM 0.19 0.33 0.07
-SeasonalForestSpring P9 ODOM 0.15 0.49 0.05
-SLAM 0.11 0.39 0.02
-Average ODOM 0.21 0.48 0.088
-SLAM 0.17 0.37 0.065
+Complete tracking results are provided in Tables 9 and 11, with all results averaged across environments.
 For the TartanAir V2 dataset, most environments utilized all available sequences.
 Sequences excluded from SLAM validation are marked with an asterisk (P*), while their corresponding visual odometry results are labeled ODOM*, indicating performance metrics computed across all available sequences for those specific environments, including the excluded sequences.
-A.1.4 Mutli-Stereo R2B Dataset
+A.1.6 Mutli-Stereo R2B Dataset
 The Multi-Stereo R2B dataset was acquired using a wheeled Carter robot platform equipped with four stereo Hawk cameras, each capturing RGB images at a resolution of 1920×1200 pixels and 30 frames per second.
@@
 Black lines represent ground truth; green lines represent odometry predictions.
-Attached table lists trajectory distances, number of 8-image batches, and trajectory numbering per environment (left to right).
+Attached table lists trajectory distances, number of 8-image batches, and trajectory index for environments (left to right).
 A.2 Appendix B: Hardware Scalability on Embedded Platforms
@@
 Consequently, for HD resolution applications, we recommend either reducing the frame rate to 15 FPS or decreasing the resolution to VGA, as our experiments revealed no significant difference in cuVSLAM visual tracking quality between VGA and HD resolutions.
-Jetson Orin Nano Super Performance.
+Jetson Orin Nano Performance.
 To evaluate the feasibility of deploying cuVSLAM on entry-level hardware, we conducted experiments replacing the Jetson Orin AGX with a Jetson Orin Nano Super operating in 25W mode and utilizing exclusively the Python stack.
@@
 The selection of appropriate resolution and frame rate parameters should be carefully balanced based on the available computational resources and specific application requirements.
-Table 8: TartanAir V2 cuVSLAM evaluation results (Envs 1–35, see Appendix A.1.3 for details)
-Env Seq Mode ATE ARE APE
+Table 9: Comparative evaluation of odometry and SLAM performance for cuVSLAM Multi-Stereo (4 stereo cameras) on the TartanGround dataset.
+To assess the impact of SLAM mode, a subset of sequences containing looped trajectories with observable loop closures was selected.
+Reported values are averages computed per environment.
+Environment Sequences Mode avgRTE avgRE RMSE APE
+Downtown P7 ODOM 0.35 0.59 0.09
+SLAM 0.25 0.41 0.09
+ForestEnv P1, P14, P17, P21 ODOM 0.42 0.78 0.16
+SLAM 0.38 0.61 0.17
+Gascola P0, P3-P6, P9 ODOM 0.17 0.41 0.08
+SLAM 0.13 0.31 0.06
+ModernCityDowntown P2, P5-P7 ODOM 0.17 0.32 0.08
+SLAM 0.12 0.24 0.06
+ModularNeighborhood P12, P20 ODOM 0.42 1.17 0.14
+SLAM 0.35 1.02 0.05
+NordicHarbor P1, P11, P17-P19 P23, P27 ODOM 0.17 0.41 0.08
+SLAM 0.10 0.26 0.04
+OldTownFall P2 ODOM 0.07 0.31 0.05
+SLAM 0.12 0.26 0.06
+OldTownSummer P2-P4 ODOM 0.09 0.29 0.06
+SLAM 0.06 0.25 0.03
+SeasonalForestAutumn P5, P8, P17 ODOM 0.20 0.32 0.07
+SLAM 0.19 0.33 0.07
+SeasonalForestSpring P9 ODOM 0.15 0.49 0.05
+SLAM 0.11 0.39 0.02
+Average ODOM 0.21 0.48 0.088
+SLAM 0.17 0.37 0.065
+Table 10: TartanAir V2 Hard cuVSLAM evaluation results (Envs 1–35, see Appendix A.1.5 for details)
+Env Seq Mode RTE RRE APE
 1 Abandoned- Cable P0-P7 ODOM 1.90 13.77 8.97
@@
 SLAM 3.70 23.22 5.60
-Table 9: TartanAir V2 cuVSLAM evaluation results (Envs 36–71, see Appendix A.1.3 for details)
-Env Seq Mode ATE ARE APE
+Table 11: TartanAir V2 Hard cuVSLAM evaluation results (Envs 36–71, see Appendix A.1.5 for details)
+Env Seq Mode RTE RRE APE
 36 ModernCity Downtown P0-P8 ODOM 2.34 11.70 4.01